### 16 Apr 2026 v2

# Cohomology of pure symmetric automorphisms of right-angled Artin groups

Abstract.
We compute the cohomology groups of the pure symmetric outer automorphism group $\Sigma\mathrm{POut}(A_{\Gamma})$ and the pure symmetric automorphism group $\Sigma\mathrm{PAut}(A_{\Gamma})$ of a right-angled Artin group $A_{\Gamma}$. Using the equivariant spectral sequence arising from the action of $\Sigma\mathrm{POut}(A_{\Gamma})$ on the generalized McCullough-Miller complex $\mathrm{MM}_{\Gamma}$, we show that $H^{q}(\Sigma\mathrm{POut}(A_{\Gamma}))$ is free abelian and we compute its rank in terms of the combinatorics of certain poset. Applying the Lyndon-Hochschild-Serre spectral sequence and the Leray-Hirsch theorem we do the same for $H^{q}(\Sigma\mathrm{PAut}(A_{\Gamma}))$. In both cases the cohomology algebra is generated in degree 1. Finally, we introduce the Generalized Brownstein-Lee Conjecture, proposing a presentation of $H^{*}(\Sigma\mathrm{PAut}(A_{\Gamma}))$, and prove that it holds in dimension $2$.
Key words and phrases:
Cohomology of pure symmetric outer automorphism groups of right angled Artin groups
2020 Mathematics Subject Classification:
Primary 20J06, 20F36; Secondary 57M07, 55P20
The author is partially supported by the Spanish Government PID2021-126254NB-I00, PID2024-155800NB-C32, and also by Universidad Pública de Navarra (grant code: Plan de promoción de grupos. “Álgebra. Aplicaciones”).
1. Introduction
Right-angled Artin groups (RAAGs) and their automorphism groups are very active field of study in geometric group theory. Given a RAAG $A_{\Gamma}$}, the automorphisms of $A_{\Gamma}$} that map each standard generator to a conjugate of itself are called pure symmetric and form a subgroup denoted $\Sigma\mathrm{PAut}(A_{\Gamma})$ that contains the subgroup of inner automorphism, so we can define the pure symmetric outer automorphism group as $\Sigma\mathrm{POut}(A_{\Gamma})\cong\Sigma\mathrm{PAut}({A_{\Gamma}})/\mathrm{Inn}({A_{\Gamma}})$. In the case when $A_{\Gamma}$ is the free group $F_{n}$, $\Sigma\mathrm{PAut}(F_{n})$ is the fundamental group of the configuration space of $n$ unknotted, unlinked circles in the sphere $S^{3}$ [9], there is also a version of this for symmetric automorphisms of other right-angled Artin groups [3].
In [15], McCullough and Miller constructed a contractible simplicial complex $\mathrm{MM}_{n}$ that admits an action of the group $\Sigma\mathrm{POut}(F_{n})$ with free abelian stabilizers. The construction is based on the Whitehead poset $\mathrm{Wh}_{n}$, whose elements can be expressed as labeled trees. In [9], Jensen, McCammond and Meier use the McCullough-Miller complex to compute $H^{*}(\Sigma\mathrm{POut}(F_{n}))$ and $H^{*}(\Sigma\mathrm{PAut}(F_{n}))$, the cohomology groups of the (outer) pure symmetric automorphisms of the free group; even more, they prove the Brownstein-Lee conjecture [5], which stablish a presentation of the cohomology ring $H^{*}(\Sigma\mathrm{PAut}(F_{n}))$.
In [1], Ardaiz-Gale and Martínez-Pérez generalized both the Whitehead poset and the construction of McCullough and Miller from the case of $\Sigma\mathrm{POut}(F_{n})$ to $\Sigma\mathrm{POut}({A_{\Gamma}})$. This new complex, $\mathrm{MM}_{\Gamma}$, admits an action of the group $\Sigma\mathrm{POut}({A_{\Gamma}})$ with free abelian stabilizers and it is also contractible. The powerful combinatorial machinery derived from the labeled trees in $\mathrm{Wh}_{n}$ is the key to prove the main results in [9]. However, vertices in $\mathrm{Wh}_{\Gamma}$ cannot be expressed as labeled trees in general, only as families of based partitions, so most arguments need to be adapted.
In this paper we will generalize some combinatorial results from $\mathrm{Wh}_{n}$ to $\mathrm{Wh}_{\Gamma}$ and use the equivariant spectral sequence derived from the $\Sigma\mathrm{POut}({A_{\Gamma}})$ action over $\mathrm{MM}_{\Gamma}$ to compute the cohomology groups $H^{q}(\Sigma\mathrm{POut}(A_{\Gamma}))$ and $H^{q}(\Sigma\mathrm{PAut}(A_{\Gamma}))$. We show (see Sections 2 and 4 for definitions):
Theorem A.
Let $K_{q}$ be the number of rank $q$ essential vertex types in $\mathrm{Wh}_{\Gamma}$, then
$H^{q}(\Sigma\mathrm{POut}(A_{\Gamma}))=\mathbb{Z}^{K_{q}}.$
Besides, the cohomology ring $H^{*}(\Sigma\mathrm{POut}(A_{\Gamma}))$ is generated by the one-dimensional classes $\bar{\gamma}_{I}^{j}$, where the $C_{I}^{j}$ are canonical automorphisms.
Theorem B.
Let $K_{i}$ be the number of rank $i$ essential vertex types in $\mathrm{Wh}_{\Gamma}$ and $N_{j}$ the number of $j$-cliques in $\Gamma$, then
$H^{q}(\Sigma\mathrm{PAut}(A_{\Gamma}))\cong\bigoplus_{i+j=q}\mathbb{Z}^{K_{i}\cdot N_{j}}.$
Besides, the cohomology ring $H^{*}(\Sigma\mathrm{PAut}(A_{\Gamma}))$ is generated in degree 1.
Furthermore, we generalize the relations from the Brownstein-Lee Conjecture to the case of RAAGs, formulate the Generalized Brownstein-Lee Conjecture, and prove that it holds in dimension 2.
Conjecture A (Generalized Brownstein-Lee Conjecture).
Let $\gamma_{A}^{i}$ be a one-dimensional class in $H^{*}(\Sigma\mathrm{PAut}(A_{\Gamma}))$, where $A$ is a connected component in $\Gamma-\mathrm{st}(v_{i})$. For each pair $v_{i},v_{j}$ of two non-adjacent vertices in $\Gamma$, let $D^{j},D^{i}$ be the respective dominant components and $\mathcal{C}$ a shared component. Then $H^{*}(\Sigma\mathrm{PAut}(A_{\Gamma}))$ is generated by the one-dimensional classes $\gamma_{A}^{i}$, modulo the following relations for each shared component $\mathcal{C}$ of $v_{i},v_{j}$.
  • (1)
    $\gamma_{A}^{i}\wedge\gamma_{A}^{i}=0$
  • (2)
    $\gamma_{A}^{i}\wedge\gamma_{B}^{j}=-\gamma_{B}^{j}\wedge\gamma_{A}^{i}$
  • (3)
    $\gamma_{D^{j}}^{i}\wedge\gamma_{D^{i}}^{j}=0$
  • (4)
    $\gamma_{\mathcal{C}}^{i}\wedge\gamma_{\mathcal{C}}^{j}=\gamma_{\mathcal{C}}^{i}\wedge\gamma_{D^{i}}^{j}+\gamma_{D^{j}}^{i}\wedge\gamma_{\mathcal{C}}^{j}$
We will denote by $R$ the ring given by the presentation in the previous conjecture.
Theorem C.
There exists a ring epimorphism
$\phi:H^{*}(\Sigma\mathrm{PAut}(A_{\Gamma}))\twoheadrightarrow R$
which is an isomorphism in degrees 1 and 2.
Another important evidence about this conjecture can be derived from the work of Martínez-Pérez and Mendonça in [14]. In Theorem A in that paper it is shown that when the graph $\Gamma$ does not contain four pairwise non-adjacent vertices $v_{1},v_{2},v_{3},v_{4}\in\Gamma$ that lie in four distinct connected components of $\Gamma-\cap_{i=1}^{4}\mathrm{lk}(v_{i})$, then the group $A_{\Gamma}$ is of Koszul type, implying that $\mathcal{U}_{\Gamma}$, the universal enveloping algebra of the rational Lie algebra associated to the descending central series of the group is quadratic and has as quadratic dual the cohomology ring of the group. It turns out [14] that for arbitrary $\Gamma$ the quadratic dual of $\mathcal{U}_{\Gamma}$ is precisely the ring with the presentation of Conjecture A. So this implies that if $\Gamma$ satisfies the condition above, then the rational version of Conjecture A holds true.
The paper is organized as follows: In Section 2 we explain some basic results about MMΓ, obtained from [1]. In Section 3, we go over the main definitions in [9]. In Section 4 we generalize the notion of one-two automorphism which plays a key role in [9] and do some technical work that will be needed later. In Section 5 we apply the previous results to the spectral sequence derived from the action of $\Sigma\mathrm{POut}(A_{\Gamma})$ on MMΓ to compute $H^{*}(\Sigma\mathrm{POut}(A_{\Gamma}))$ and prove Theorem A. In Section 6, we use the Lyndon-Hoschild-Serre spectral sequence associated to $1\to A_{\Gamma}\to\Sigma\mathrm{PAut}(A_{\Gamma})\to\Sigma\mathrm{POut}(A_{\Gamma})\to 1$ and the Leray-Hirsch Theorem to compute the $H^{*}(\Sigma\mathrm{PAut}(A_{\Gamma}))$ groups and prove Theorem B. In Section 7 we take the first steps into proving the Brownstein-Lee conjecture for $\Sigma\mathrm{PAut}(A_{\Gamma})$ and we prove Theorem C.
Acknowledgments: The author would like to thank Ric Wade for helpful conversations during a stay in Oxford that made spectral sequences less ghostly, and Conchita Martínez-Pérez for her invaluable help.
2. The McCullough-Miller complex
In this chapter we will explain some basic results about the McCullough-Miller complex for RAAGs. A more detailed explanation can be found at [1].
Definition 2.1 (Graph).
We will use $\Gamma$ to denote a finite simplicial graph with vertex set $V(\Gamma)$ and set of edges $E(\Gamma)$. Very often we will identify $\Gamma$ or a subgraph of $\Gamma$ with its vertex set. Let $v\in V(\Gamma)$, $\mathrm{lk}(v)$ denotes the link of $v$, which is the subgraph induced by the vertices adjacent to $v$. We will define the star of $v$, $\mathrm{st}(v)$, as the subgraph induced by $\mathrm{lk}(v)\cup\{v\}$.
Associated to $\Gamma$ we may define a group as follows.
Definition 2.2 (RAAG).
The RAAG (right-angled Artin group) defined by a simplicial graph $\Gamma$ is the group with presentation
$A_{\Gamma}=\langle v\in V(\Gamma)\mid[v,w]=1\textrm{ whenever }\{v,w\}\in E(\Gamma)\rangle$
This is, $A_{\Gamma}$} is the group generated by the vertices of $\Gamma$, which are called standard generators so that two standard generators commute if and only if they are adjacent in $\Gamma$. Two simple examples are the extreme cases: if $\Gamma$ is a graph with no edges, then $A_{\Gamma}$ is a free group, and if $\Gamma$ a complete graph, then $A_{\Gamma}$ is a free abelian group.
2.1. Pure symmetric automorphisms of RAAGs
Definition 2.3 (Pure symmetric (outer)automorphisms).
Let $A_{\Gamma}$ be a RAAG. The group of pure symmetric automorphisms $\Sigma\mathrm{PAut}(A_{\Gamma})$ of $A_{\Gamma}$ is the group consisting of those automorphisms of $A_{\Gamma}$ that map each standard generator (i.e., each vertex of $\Gamma$) to a conjugate of itself. Obviously, it contains the inner automorphism and the corresponding external version is denoted $\Sigma\mathrm{POut}(A_{\Gamma})=\Sigma\mathrm{PAut}(A_{\Gamma})/\mathrm{Inn}(A_{\Gamma}).$
Next, we introduce an important family of elements of $\Sigma\mathrm{PAut}(A_{\Gamma})$.
Definition 2.4 (Partial conjugations).
Given $v\in L=V(\Gamma)\cup V(\Gamma)^{-1}$ and $A$ a union of connected components of $\Gamma-\mathrm{st}(v)$, the partial conjugation of $A$ by $v$ is the automorphism $C_{A}^{v}$ given by
$\left\{\begin{array}[]{l}C_{A}^{v}(w)=vwv^{-1},\text{ if }w\in A\\ C_{A}^{v}(z)=z,\text{ if }z\notin A\end{array}\right$.
where $w,z\in\Gamma$.
Theorem 2.5 (Laurence [12]).
The set of partial conjugations is a generating system for $\Sigma\mathrm{PAut}(A_{\Gamma})$.
Remark.
For the rest of the paper, we will only consider RAAGs $A_{\Gamma}$} such that $\Gamma$ has no dominating vertex. Equivalently, RAAGs with trivial center. There is no loss of generality in this assumption, which will be helpful in Section 6, because, by Laurence’s Theorem, $\Sigma\mathrm{PAut}({A_{\Gamma}})\cong\Sigma\mathrm{PAut}(A_{\Gamma^{\prime}})$ if $\Gamma$ is obtained from $\Gamma^{\prime}$ by removing all the dominating vertices.
When we are working with non adjacent vertices $u$ and $v$, it will be really useful to think about $\Gamma-\mathrm{st}(u)$ in terms of graph components. We will use the next notation due to Day and Wade [7].
Definition 2.6 (Graph components: shared, dominant and subordinate).
Let $u,v\in\Gamma$ be not linked. A connected component of $\Gamma-\mathrm{st}(u)$ that is also a connected component of $\Gamma-\mathrm{st}(v)$ is said to be shared. The unique component of $\Gamma-\mathrm{st}(u)$ that contains $v$ is called the dominant component of $\Gamma-\mathrm{st}(u)$ respect to $v$, we will denote it by $D^{v}$ when $u$ is clear. Finally, a subordinate component is any connected component of $\Gamma-\mathrm{st}(u)$ that is contained in the dominant component of $\Gamma-\mathrm{st}(v)$ respect to $u$. By Lemma 2.1 in [7], any connected component of $\Gamma-\mathrm{st}(u)$ is of one of these types. Observe that if $A$ is a connected component of $\Gamma-\mathrm{st}(u)$ which is either shared or subordinate, then $A\cap\mathrm{st}(v)=\emptyset$.
Using the terminology in the previous definition, one can rewrite the presentation of Koban and Piggott ([10]) as follows:
Theorem 2.7.
The group $\Sigma\mathrm{PAut}(A_{\Gamma})$ is the group generated by all partial conjugations $C_{A}^{u}$ for $u\in\Gamma$ and $A$ a connected component of $\Gamma-\mathrm{st}(u)$, subject to the following relations:
  • i)
    $\left[C_{A}^{u},C_{B}^{v}\right]=1$ if $u\in\mathrm{st}(v)$,
  • ii)
    $\left[C_{A}^{u},C_{B}^{v}\right]=1$ if $u\notin\mathrm{st}(v)$ and either $A$ and $B$ are shared but distinct or at least one of $A,B$ is subordinate,
  • iii)
    $\left[C_{A}^{u}C_{B}^{u},C_{A}^{v}\right]=1$ if $u\notin\mathrm{st}(v)$, $A$ is shared and $B$ dominant.
Definition 2.8 (Valid based partitions, petals and length).
Let $u\in\Gamma$ be a vertex. A $\Gamma$-valid based partition with operative factor $u$ is a partition of the form $\tau_{u}=\{\{u\},P_{1},\ldots,P_{k}\}$ of the set $\Gamma-\mathrm{lk}(u)$ so that each $P_{i}$ is a union of connected components. Each $P_{i}$ is called a petal of the based partition. The number of petals, $k$, is the length of the based partition, and we denote it $l(\tau_{u})$.
From now on, every based partition we use will be a $\Gamma$-admissible based partition, for simplicity we will simply say based partition if the graph $\Gamma$ is clear.
Definition 2.9 (Crossings, disjoint partition and compatibility).
We say that two based partitions $\tau_{u}$ and $\tau_{v}$ cross if $u\notin\mathrm{st}(v)$ and there are petals $P$ of $\tau_{u}$ and $Q$ of $\tau_{v}$ such that $D^{v}\not\subset P$, $D^{u}\not\subset Q$, and $P\cap Q\neq\emptyset$. It is easy to check that this is equivalent to asking that there is at least one shared component $\mathcal{C}$ such that $\mathcal{C}\subseteq P$ and $\mathcal{C}\subseteq Q$. We will say that two based partitions $\tau_{u},\tau_{v}$ are disjoint if $u\notin\mathrm{st}(v)$ and they do not cross and we say that they are compatible if they are disjoint or $u\in\mathrm{st}(v)$.
Definition 2.10 (Vertex type).
We define a vertex type $\tau=\{\tau_{v_{1}},...,\tau_{v_{n}}\}$ as a collection of pairwise compatible based partitions, such that there is one for each vertex in $\Gamma$.
We will say that two vertex types $\tau$ and $\tau^{\prime}$ are compatible if their based partitions are pairwise compatible. And we will say that a collection of vertex types is compatible if they are pairwise compatible.
Remark.
Here we are using a slightly different notation from [1], where vertex types where denoted $\underline{\underline{A}}$, and a based partition $\underline{A}_{u}$. The extensive notation in [1] is forced by the technical work done there. Here, this simpler version will be enough.
Definition 2.11 (Order).
Let $v\in\Gamma$ and $\tau_{v}$, $\tau^{\prime}_{v}$ two based partitions with operative factor $v$. We say that $\tau_{v}\leq\tau_{v}^{\prime}$ if every petal of $\tau_{v}^{\prime}$ is contained in a petal of $\tau_{v}$, i.e., when every petal of $\tau_{v}$ is a union of petals of $\tau_{v}^{\prime}$. In particular, if $\tau_{v}\leq\tau_{v}^{\prime}$, then $l(\tau_{v})\leq l(\tau^{\prime}_{v})$. For vertex types, we will say that $\tau\leq\tau^{\prime}$ if $\tau_{v}\leq\tau^{\prime}_{v},\ \textrm{ for every }v\in V(\Gamma)$.
We have the next Lemma from [1, Lemma 4.5.]:
Lemma 2.12.
Let $\tau$ and $\tau^{\prime}$ be vertex types with $\tau<\tau^{\prime}$, then $\tau$ and $\tau^{\prime}$ are compatible.
Definition 2.13 ($\Gamma$-Whitehead poset).
The $\Gamma$-Whitehead poset WhΓ is the poset formed by all the vertex types associated to $\Gamma$ under the partial order previously defined.
Definition 2.14 (Nuclear vertex).
The nuclear vertex type, denoted $\mathcal{O}$, is the vertex type such that for each $v\in\Gamma$ the associated based partition is trivial, i.e., $\mathcal{O}_{v}=\{\{v\},P\}$, where $P=\Gamma-\mathrm{st}(v)$. It is the unique minimal element in the $\mathrm{Wh}_{\Gamma}$ poset.
Definition 2.15 (Rank of a vertex type).
We denote by $l(\tau_{v})$ the length of a based partition, i.e., the number of petals it has. Then, the rank of a vertex type $\tau$ is
$\mathrm{rk}(\tau)=\displaystyle\sum_{v\in\Gamma}(l(\tau_{v})-1)$
Definition 2.16 (Carried automorphism).
The group of automorphisms carried by a based partition $\tau_{u}$ is the subgroup of $\Sigma\mathrm{PAut}({A_{\Gamma}})$ generated by the partial conjugations $C^{u}_{I}$ where $I$ runs over the petals of $\tau_{u}$. The group of automorphisms carried by a vertex type $\tau$ is the subgroup of $\Sigma\mathrm{PAut}({A_{\Gamma}})$ generated by the automorphisms carried by each based partition $\tau_{u}\in\tau$.
The following is an easy observation that will be useful later.
Lemma 2.17.
Let $\tau\in\mathrm{Wh}_{\Gamma}$ be a vertex type. The subgroup of $\Sigma\mathrm{PAut}({A_{\Gamma}})$ carried by $\tau$ contains the inner automorphisms. Moreover, given two vertex types $\tau$ and $\tau^{\prime}$ such that $\tau<\tau^{\prime}$, the subgroup of automorphisms carried by $\tau$ is inside the subgroup of automorphisms carried by $\tau^{\prime}$.
In what follows, we will often consider the subgroups of outer automorphisms carried by either a given based partition or a vertex type, which are just the corresponding quotients over $\mathrm{Inn}(A_{\Gamma})$.
Definition 2.18 (Marked vertex type and the action of $\Sigma\mathrm{PAut}(A_{\Gamma})$).
We say that a generating system $X$ of $A_{\Gamma}$ is a basis if there is some $\alpha\in\Sigma\mathrm{PAut}(A_{\Gamma})$ such that $X=\alpha(V(\Gamma))$. A marked based partition is a pair $(X,\tau_{u})$ where $X$ is an ordered basis of $A_{\Gamma}$ and $\tau_{u}$ is a based partition and a marked vertex type $(X,\tau)$ is the family of marked based partitions $(X,\tau_{u})$ where $\tau_{u}$ runs over the set of based partitions of a vertex type $\tau$. By marking a based partition $\tau_{u}$ we understand that we are relabeling the elements in the partition using $X$ (via the automorphism $\alpha$). We extend the partial order of Definition 2.11 to marked vertex types as follows: $(X_{1},\tau)\leq(X_{2},\tau^{\prime})$ if $X_{1}=X_{2}$ and $\tau\leq\tau^{\prime}$. The marked vertex types form a poset and the group $\Sigma\mathrm{PAut}(A_{\Gamma})$ acts on this poset via
$\alpha(X,\tau)=(\alpha(X),\tau)$
where $\alpha\in\Sigma\mathrm{PAut}(A_{\Gamma})$. We say that $\alpha$ is carried by a marked vertex type $(X,\tau)$ if $\gamma\alpha\gamma^{-1}$ is carried by $\tau$ where $\gamma(X)=V(\Gamma)$.
Next, we define an equivalence relationship in the set of marked vertex types as follows.
Definition 2.19.
We say that two marked vertex types $(X_{1},\tau)$ and $(X_{2},\tau)$ are equivalent and we put $(X_{1},\tau)\equiv(X_{2},\tau)$ if there exists an $\alpha\in\Sigma\mathrm{PAut}(A_{\Gamma})$ carried by $(X_{1},\tau)$ such that $\alpha(X_{1})=X_{2}$. We will denote by $X_{1},\tau]$ the equivalence class of $(X_{1},\tau)$ under this relation.
Definition 2.20 ($\Gamma$-McCullough-Miller poset).
The $\Gamma$-McCullough-Miller poset is the poset formed by the set of equivalence classes under the equivalence relationship $\equiv$ defined above with the induced partial order. The group $\Sigma\mathrm{PAut}(A_{\Gamma})$ acts on this poset via
$\alpha([X,\tau])=[\alpha(X),\tau]$
and this action factors through the epimorphism $\Sigma\mathrm{PAut}(A_{\Gamma})\to\Sigma\mathrm{POut}(A_{\Gamma})$ so we also have an action of $\Sigma\mathrm{POut}(A_{\Gamma})$.
The $\Gamma$-McCullough-Miller poset is formed by copies of the $\Gamma$-Whitehead poset attached to each other through the equivalence relation $\equiv$. It is easy to verify that nuclear vertex types only carry inner automorphisms, so there is a bijection between the set of nuclear vertices $X,\mathcal{O}]$ in the $\Gamma$-McCullough-Miller poset and the set of conjugacy classes of bases of $A_{\Gamma}$.
Definition 2.21 ($\mathrm{MM}_{\Gamma}$ complex).
We define the McCullough-Miller complex $\mathrm{MM}_{\Gamma}$ for a RAAG $A_{\Gamma}$ as the simplicial realization of the $\Gamma$-McCullough-Miller poset. In the case when $A_{\Gamma}$} is the free group with rank $n$ we will denote this complex by $\mathrm{MM}_{n}$.
Remark.
The definition of the action above implies that the stabilizer $H=\mathrm{Stab}_{\Sigma\mathrm{POut}({A_{\Gamma}})}[V(\Gamma),\tau]$ of the vertex $V(\Gamma),\tau]$ under the $\Sigma\mathrm{POut}({A_{\Gamma}})$ action is precisely the subgroup generated by those automorphisms carried by the vertex type $\tau$. This implies that the stabilizer of the vertex $X,\tau]$ is $\alpha H\alpha^{-1}$ where $\alpha$ is such that $X=\alpha(V(\Gamma))$. It is proved in [1, Theorem A] that these stabilizers are free abelian, that their rank can be computed in terms of the length of the based partitions of $\tau$ and that they yield in fact all the cell stabilizers of the McCullough-Miller complex $\mathrm{MM}_{\Gamma}$.
Taking this into account, one sees that the McCullough-Miller complex $\mathrm{MM}_{\Gamma}$ can be also defined as follows. Recall that for a poset $\mathcal{F}$ of subgroups of a group $G$, one can define a new poset, called the coset poset, as
$C(\mathcal{F})=\{gH\mid H\in\mathcal{F},g\in G\}.$
Then, the McCullough-Miller poset is the coset poset of the family
$\{\mathrm{Stab}_{\Sigma\mathrm{POut}(A_{\Gamma})}[V(\Gamma),\tau]\mid\tau\text{ vertex type}\}$
of subgroups of $\Sigma\mathrm{POut}(A_{\Gamma})$ and $\mathrm{MM}_{\Gamma}$ is the associated geometric realization.
[…]
3. The one-two basis
Our main goal is to compute $H^{*}(\Sigma\mathrm{POut}(A_{\Gamma}))$, and we will do it using an equivariant spectral sequence for the $\Sigma\mathrm{POut}({A_{\Gamma}})$ action on $\mathrm{MM}_{\Gamma}$. If a group acts simplicially on a contractible complex, then the first page of this equivariant spectral sequence can be computed in terms of the cell stabilizers (Chapter VII, Proposition 7.3 in [4]). A key step in this process is choosing a good basis for these stabilizers. We will do it in a way that generalizes the choice of [9].
As proved in [1] (Lemma 5.3.) and explained in the Remark after Definition 2.21, the stabilizer of a vertex $\tau$ is precisely the group of automorphisms carried by $\tau$. By Definition 2.16, the partial conjugations $C_{I}^{u}$, where $I$ runs over the petals of the based partitions $\tau_{u}$ of $\tau$, are a generating set for the stabilizer. Observe that for a based partition $\tau_{u}=\{\{u\},I_{1},...,I_{k}\}$, the composition $C_{I_{1}}^{u}\cdots C_{I_{k}}^{u}$ is the inner automorphism associated to $u$. Since we are working with outer-automorphisms, this means that the set $\{C_{I_{1}}^{u},\cdots,C_{I_{k}}^{u}\}$ is not independent. So, if we want a minimal generating set, we have to choose which one to leave aside. This choice problem is solved as follows in [9] for the case of the pure symmetric outer-automorphisms of free group:
Definition 3.1 (One-two automorphism).
Let $\{x_{1},...,x_{n}\}$ be a fixed basis for the free group of rank $n$ $F_{n}$. A one-two automorphism is a partial conjugation $C_{I}^{x_{j}}$ such that:
  1. 1)
    $x_{1}\notin I$.
  2. 2)
    if $x_{j}=x_{1}$, then $x_{2}\notin I$.
In simple words, a one-two automorphism never conjugates $x_{1}$ and never conjugates $x_{2}$ by $x_{1}$. The last important definition that we need from [9] is the notion of essential vertex type.
Remark.
In $\mathrm{Wh}_{n}$, a vertex type $\tau$ can be expressed as a labeled tree, and the order in the Whitehead poset can be defined by folding or unfolding these trees. When a tree is folded (or unfolded) we are just joining (or splitting) petals in the based partitions. The nice property about these labeled trees is that they only allow precisely those splits that will result in another vertex type. To generalize this idea, if $\tau$ is a vertex type in $\mathrm{Wh}_{\Gamma}$ for an arbitrary $\Gamma$, we say that a petal $I$ in a based partition $\tau_{u}$ of $\tau$ can be split when $\tau^{\prime}_{u}$, the based partition that results from the split, is still compatible with every other based partition $\tau_{v}$ of $\tau$.
Definition 3.2.
A vertex $\tau$ in $\mathrm{Wh}_{n}$ is essential if for each based partition $\tau_{j}$ of $\tau$ we have that:
  • 1)
    If $x_{j}=x_{1}$, the only petal that can be split to yield a new vertex type is the one containing $x_{2}$.
  • 2)
    If $x_{j}\neq x_{1}$, the only petal that can be split to yield a new vertex type is the one containing $x_{1}$.
4. Choosing the basis for $\Sigma\mathrm{POut}({A_{\Gamma}})$
From now on we fix an order in the elements in $V(\Gamma)$, and denote them by $v_{1},...,v_{n}$. Sometimes, for simplicity and in the aim of a clear notation, we will just identify $v_{i}$ with $i$; in this sense, we will often put $\tau_{i}$ to denote the based partition $\tau_{v_{i}}$, or $C^{j}_{I}$ for the partial conjugation $C^{v_{j}}_{I}$. Let $\tau$ be a vertex type in MMΓ and $\tau_{i}$ a based partition of $\tau$. The product of the partial conjugations associated to each petal of $\tau_{i}$ gives the inner automorphism associated to $v_{i}$, so it is trivial in the outer group. We need to have a way to choose families of partial conjugations associated to $\tau_{i}$ which are independent.
Definition 4.1 (Canonical automorphism).
Let $\tau$ be a vertex type in WhΓ and $\tau_{j}$ a based partition of $\tau$. For each petal $I$ in $\tau_{j}$ is the associated partial conjugation is $C_{I}^{j}$. Let $k$ be the smallest label in $\Gamma-\mathrm{st}(v_{j})$. We say that $C_{I}^{j}$ is a canonical automorphism if $v_{k}\notin I$. The image of $C_{I}^{j}$ in $\Sigma\mathrm{POut}(A_{\Gamma})$, which we will denote by $\bar{C}_{I}^{j}$, is called a canonical outer-automorphism.
It is easy to check that, in the case of the free group, these are precisely the one-two automorphisms in Definition 3.1.
Example 4.2.
Let $\Gamma$ be as in Figure 1, and
$\tau_{2}=\{\{2\},\{4\},\{5,6,7,8,9,10,11\}\}$
the only non-trivial based partition associated to $v_{2}\in\Gamma$. The smallest label in $\Gamma-\mathrm{st}(v_{2})$ is 4, so the only canonical automorphism associated to $v_{2}$ is the partial conjugation $C_{I}^{2}$ where $I=\{5,6,7,8,9,10,11\}$.
2345689710111
Figure 1. A graph $\Gamma$ for Example 4.2
Definition 4.3 (Canonical basis).
[…]
  • 1)
    II is a petal in a based partition $\tau_{j}$ of $\tau$.
  • 2)
    $\bar{C}^{j}_{I}$ is a canonical outer-automorphism.
We denote it by $\mathcal{B}(\tau)$. If $\sigma$ is a simplex in $|\mathrm{Wh}_{\Gamma}|$ corresponding to a chain $\tau^{0}<\cdots<\tau^{k}$ then the stabilizer of $\sigma$ is the stabilizer of $\tau^{0}$ so we define $\mathcal{B}(\sigma):=\mathcal{B}(\tau^{0})$.
Following the explanation in the Remark after Definition 2.21, we have that
Lemma 4.4.
[…]
Definition 4.5 (Support).
Let $\mathcal{A}$ be a family of canonical outer-automorphisms, then we define its support as
$\textrm{supp}(\mathcal{A}):=\{\sigma\in|\mathrm{Wh}_{\Gamma}|\ |\ \mathcal{A}\subseteq\mathcal{B}(\sigma)\}.$
Remark.
$\mathrm{supp}(\mathcal{A})$ is not necessarily a subcomplex of $|\mathrm{Wh}_{\Gamma}|$.
Definition 4.6 (Compatibility).
A collection $\mathcal{A}$ of canonical outer-automorphisms is called compatible when $\mathrm{supp}(\mathcal{A})\neq\emptyset$.
Example 4.7.
[…]
We have that $\bar{C}_{\{4,5\}}^{3}$ is a canonical outer-automorphism since $v_{1}$ does not get conjugated. We note that $\bar{C}_{\{4,5\}}^{3}\in\mathcal{B}(\tau)$ since it is canonical and $\{4,5\}$ corresponds to a petal in the based partition $\tau_{3}$ of $\tau$; in fact, $\mathcal{B}(\tau)=\{\bar{C}_{\{4,5\}}^{3}\}$. Now, $\tau^{\prime}$ has been obtained from $\tau$ by splitting the petal $\{4,5\}$, so we have that $\mathcal{B}(\tau^{\prime})=\{\bar{C}_{\{4\}}^{3},\bar{C}_{\{5\}}^{3}\}$. It is obvious that $\bar{C}_{\{4,5\}}^{3}\in\mathrm{Stab}(\tau^{\prime})$, since it is the product of the two elements in the basis. However, $\bar{C}_{\{4,5\}}^{3}\notin\mathcal{B}(\tau^{\prime})$, since $I=\{4,5\}$ is not associated to a single petal in the based partition $\tau^{\prime}_{3}$ of $\tau^{\prime}$.
This also gives us an example of $\mathrm{supp}(\mathcal{A})$ not being a subcomplex in $|\mathrm{Wh}_{\Gamma}|$. Let $\mathcal{A}=\{\bar{C}_{\{4,5\}}^{3}\}$; if we take the simplex $\sigma$ given by the chain $\tau<\tau^{\prime}$, we have that $\sigma\in\mathrm{supp}(\mathcal{A})$ since $\mathcal{B}(\sigma)=\mathcal{B}(\tau)$. However, $\tau^{\prime}\notin\mathrm{supp}(\mathcal{A})$, so it is not a subcomplex.
[…]
Let $\mathcal{A}$ be a family of canonical outer-automorphisms, then we define the following complexes
$\mathrm{Peripheral}(\mathcal{A})=\{\sigma\in|\mathrm{Wh}_{\Gamma}|\ |\ \mathcal{A}\subseteq\mathrm{Stab}(\sigma),\mathcal{A}\nsubseteq\mathcal{B}(\sigma)\}$
and
$\mathcal{C}(\mathcal{A})=\{\sigma\in|\mathrm{Wh}_{\Gamma}|\ |\ \mathcal{A}\subseteq\mathrm{Stab}(\sigma)\}.$
Example 4.9.
[…]
Definition 4.10 (Cone point of $\mathcal{A}$).
For any compatible set $\mathcal{A}$ there is a unique vertex type $\tau$ such that $\mathcal{B}(\tau)=\mathcal{A}$. This vertex type is called the cone point of $\mathcal{A}$ and denoted by $\tau(\mathcal{A})$.
Following the definition, we have that if $\bar{C}_{I_{1}}^{i},...,\bar{C}_{I_{k}}^{i}$ is the set of canonical outer-automorphisms of $\mathcal{A}$ that conjugate by $v_{i}$, then the based partition associated to $v_{i}$ in $\tau(\mathcal{A})$ will be $\tau(\mathcal{A})_{i}=\{\{v_{i}\},I_{1},...,I_{k},M\}$; where $M=\Gamma-(\mathrm{st}(v_{i})\cup I_{1}\cup\cdots\cup I_{k})$, and $M$ contains the minimal element of $\Gamma-\mathrm{st}(v_{i})$. It is easy to check that for all vertex types $\tau$ the cone point of its canonical basis is $\tau$ itself; i.e., $\tau(\mathcal{B}(\tau))=\tau$. And for all compatible collections $\mathcal{A}$, the canonical basis of its cone point is $\mathcal{A}$ itself; i.e., $\mathcal{B}(\tau(\mathcal{A}))=\mathcal{A}$. As a consequence, the cone point $\tau(\mathcal{A})$ lies in $\mathrm{supp}(\mathcal{A})$.
The next Lemma is a direct consequence of Lemma 2.17.
Lemma 4.11.
[…]
Lemma 4.12.
Both $\mathrm{Peripheral}(\mathcal{A})$ and $\mathcal{C}(\mathcal{A})$ are subcomplexes of $\mathrm{Wh}_{\Gamma}$.
Proof.
Let $\sigma\in\mathrm{Peripheral}(\mathcal{A})$ with $\sigma=\tau^{0}<\cdots<\tau^{k}$. We have to check that every vertex type $\tau^{i}$ is also in $\mathrm{Peripheral}(\mathcal{A})$. By definition of $\mathrm{Peripheral}(\mathcal{A})$, we have that $\mathcal{A}\subseteq\mathcal{B}(\sigma)$ and $\mathcal{A}\nsubseteq\mathcal{B}(\sigma)$, which means that $\mathcal{A}\subseteq\mathrm{Stab}(\tau^{0})$ and $\mathcal{A}\nsubseteq\mathcal{B}(\tau^{0})$. We have to check that $\mathcal{A}\subseteq\mathrm{Stab}(\tau^{i})$ and $\mathcal{A}\nsubseteq\mathcal{B}(\tau^{i})$ for every $\tau^{i}$. Since $\mathcal{A}\subseteq\mathrm{Stab}(\tau^{0})$ and $\tau^{0}<\tau^{i}$, by Lemma 2.17, we have that $\mathcal{A}\subseteq\mathrm{Stab}(\tau^{i})$.
Now, since $\mathcal{A}\nsubseteq\mathcal{B}(\tau^{0})$, there must be a canonical outer-automorphism $\bar{C}^{j}_{I}\in\mathcal{A}$ such that $\bar{C}^{j}_{I}=\bar{C}^{j}_{I_{1}}\cdots\bar{C}^{j}_{I_{k}}$, where $k>1$ and all the $\bar{C}^{j}_{I_{i}}\in\mathcal{B}(\tau^{0})$, i.e, all the $I_{i}$ are petals in the based partition $\tau^{0}_{j}$ of $\tau^{0}$. As $\tau^{0}<\tau^{i}$, all the petals in the based partition $\tau^{i}_{j}$ are either the same or a result of splitting some petals in $\tau^{0}_{j}$. As a consequence, we have that $\bar{C}^{j}_{I}\notin\mathcal{B}(\tau^{i})$, because in other case $I=I_{1}\cup\cdots\cup I_{k}$ would be a petal in the based partition $\tau^{i}_{j}$ of $\tau^{i}$, which is a contradiction.
Applying Lemma 2.17, it is obvious that $\mathcal{C}(\mathcal{A})$ is a subcomplex. ∎
Lemma 4.13.
Let $\mathcal{A}$ be a family of canonical outer-automorphisms, then $\mathcal{C}(\mathcal{A})=\{\sigma\in|\mathrm{Wh}_{\Gamma}|\ |\ \tau(\mathcal{A})\in\sigma\}$ and, as a consequence, $\mathcal{C}(\mathcal{A})$ is contractible (is a cone over $\tau(\mathcal{A})$).
Proof.
It follows from Lemma 4.11 and the definition of $\mathcal{C}(\mathcal{A})$. ∎
Recall that, as explained in the Remark after Definition 3.1, for a vertex type $\tau$, we cannot split petals arbitrarily if we want to obtain another vertex type. A petal $I$ in a based partition $\tau_{i}$ of $\tau$ might not admit any split because of two possible reasons:
  • a)
    The petal is formed by a single connected component in $\Gamma-\mathrm{st}(v_{i})$, so it cannot be divided anymore.
  • b)
    There is another based partition $\tau_{j}$ in $\tau$, such that $v_{i},v_{j}$ have a shared component; and the splitting of the petal in $\tau_{i}$ containing the dominant component $D^{j}$ produces a crossing with $\tau_{j}$, breaking the compatibility.
We are now ready to generalize the idea of essential vertices:
Definition 4.14 (Essential).
[…]
Let us explain the previous definition: the canonical basis is defined using the partial conjugations associated to the petals that do not contain the minimal element. If we split a petal which does not contain the minimal element, then the canonical basis will change. If we have an essential collection $\mathcal{A}$, then for each $\tau^{\prime}$ such that $\tau(\mathcal{A})\leq\tau^{\prime}$ we have that $\tau^{\prime}$ supports $\mathcal{A}$, since the petals that defined the canonical outer-automorphisms in $\mathcal{A}$ cannot be divided. So, in a simple way, having an essential collection $\mathcal{A}$ ensures that if $\tau$ supports $\mathcal{A}$ and $\tau\leq\tau^{\prime}$, then $\tau^{\prime}$ also supports $\mathcal{A}$.
Lemma 4.15.
Let $\mathcal{A}$ be a collection of compatible canonical outer-automorphisms. Then $\mathcal{A}$ is essential if and only if the complex $\mathrm{Peripheral}(\mathcal{A})$ is empty and if and only if $\mathrm{supp}(\mathcal{A})$ is a subcomplex of $|\mathrm{Wh}_{\Gamma}|$. In this case, moreover, $\mathcal{C}(\mathcal{A})=\mathrm{supp}(\mathcal{A})$ so $\mathrm{supp}(\mathcal{A})$ is contractible.
Proof.
Let $\mathcal{A}$ be essential, $\tau=\tau(\mathcal{A})$ and $\bar{C}_{I}^{j}\in\mathcal{A}$. By definition of canonical and essential, the petal $I$ in the based partition $\tau_{j}$ does not contain the minimal element, so it cannot be divided. Since every vertex $\tau^{\prime}$ with $\tau<\tau^{\prime}$ is obtained by splitting petals in $\tau$, we have that in the based partition $\tau^{\prime}_{j}$ we will still have the petal $I$ (since it cannot be divided), so $\bar{C}_{I}^{j}\in\mathcal{B}(\tau^{\prime})$ and $\mathcal{A}\subset\mathcal{B}(\tau^{\prime})$.
[…]
For the second equivalence, as explained in the previous paragraph, $\mathcal{A}$ is essential if and only if when $\tau$ supports $\mathcal{A}$ and $\tau\leq\tau^{\prime}$, then $\tau^{\prime}$ also supports $\mathcal{A}$. As a consequence, if $\sigma=\tau^{0}<\cdots<\tau^{k}\in\mathrm{supp}(\mathcal{A})$, then for any $\tau^{i}$ we have that $\tau^{i}\in\mathrm{supp}(\mathcal{A})$. ∎
We now turn to the case when $\mathcal{A}$ is inessential. It makes sense to think that if we take an inessential vertex type and we keep splitting its petals, at some point we should get an essential vertex type. We aim to do this process in an ordered way.
Definition 4.16 (Worrisome petal).
Let $\tau$ be an inessential vertex type. We call a petal in a based partition $\tau_{i}$ of $\tau$ worrisome if it can be split and it does not contain the minimal element in $\Gamma-\mathrm{st}(v_{i})$.
Remark.
Observe that if $\tau$ is inessential, there must be at least one worrisome petal; and if it is essential there is none. Further, we know that if $\tau=\tau(\mathcal{A})$ is inessential, then $\mathrm{Peripheral}(\mathcal{A})$ is not empty, and now we have another characterization: let $\tau^{\prime}$ be a vertex type, then $\tau^{\prime}\in\mathrm{Peripheral}(\mathcal{A})$ if and only if $\tau^{\prime}>\tau$ and at least one of the worrisome petals in $\tau$ is split in $\tau^{\prime}$.
We will state some technical lemmas about worrisome petals:
Lemma 4.17.
Let $\tau$ be an inessential vertex type and $\tau_{i}=\{\{v_{i}\},P_{1},...,P_{k}\}$, $\tau_{j}=\{\{v_{j}\},Q_{1},...,Q_{m}\}$ two based partitions of $\tau$ such that $v_{i}$ and $v_{j}$ are not adjacent. Let us suppose that $P_{1}$ contains $D^{j}$ and $Q_{1}$ contains $D^{i}$, the respective dominant components. Then, $P_{1}$ and $Q_{1}$ cannot both be worrisome.
Proof.
We have two different cases:
  • (i)
    The minimal elements in $\Gamma-\mathrm{st}(v_{i})$ and $\Gamma-\mathrm{st}(v_{j})$ are the same: In this case the minimal element is either in the dominant component of one of them (and in a subordinate component of the other) or in a shared component. If it is in the dominant of one of them, say $D^{j}$, then $P_{1}$ is not worrisome since it contains the minimal element and we have finished. If it is in a shared component, we have that the shared component must lie in the same petal as the dominant component in at least one of the two based partitions, if not a cross would occur between them, contradicting the fact that the based partitions are compatible; so that petal is not worrisome.
  • (ii)
    The minimal element in $\Gamma-\mathrm{st}(v_{i})$ is different from the one in $\Gamma-\mathrm{st}(v_{j})$: We can assume that the smallest one is the one in $\Gamma-\mathrm{st}(v_{i})$. This means that the vertex with the smallest label in $\Gamma-\mathrm{st}(v_{i})$ is in $\mathrm{st}(v_{j})$, as a consequence, it is in $D^{j}$, so $P_{1}$ is not worrisome.
Lemma 4.18.
Let $\tau$ be an inessential vertex type and $\tau_{i}=\{\{v_{i}\},I,P_{1},...,P_{k}\}$, $\tau_{j}=\{\{v_{j}\},J,Q_{1},...,Q_{m}\}$ two based partitions of $\tau$, such that both $I,J$ are worrisome petals. Let $\tau^{\prime}_{i}=\{\{v_{i}\},I_{1},I_{2},P_{1},...,P_{k}\}$ and $\tau^{\prime}_{j}=\{\{v_{j}\},J_{1},J_{2},Q_{1},...,Q_{m}\}$ be based partitions obtained by splitting the worrisome petals in an admissible way. Then $\tau^{\prime}_{i}$ and $\tau^{\prime}_{j}$ are compatible.
Proof.
The case when $v_{i}$ and $v_{j}$ are adjacent is trivial, since based partitions associated to adjacent vertices are always compatible. We study the case when they are not adjacent. Since $\tau^{\prime}_{i}$ is obtained by an admissible split, it means that it is compatible with al the based partitions of $\tau$, in particular, it is compatible with $\tau_{j}$. By the same reason, $\tau^{\prime}_{j}$ is compatible with $\tau_{i}$. By Lemma 4.17, we have that at least one of the worrisome petals does not contain the dominant component.
By symmetry we assume that $I$ is the one not containing $D^{j}$. Let us suppose that $\tau^{\prime}_{i}$ is not compatible with $\tau^{\prime}_{j}$, i.e., there is crossing between them. By Lemma 5.11 in [1], if $\tau_{i}$ is compatible with $\tau^{\prime}_{j}$ but $\tau^{\prime}_{i}$ is not, then the petal containing the dominant component in $\tau_{i}$ is split, but this is a contradiction, since the only petal that we have split is $I$, and $I$ does not contain the dominant component.
In general, for a vertex type $\tau$, splitting a petal in a based partition $\tau_{i}$ of $\tau$ might imply that we cannot split some other petal in another based partition $\tau_{j}$ of $\tau$, because crossings might appear. The previous lemma gives us a really important result for the rest of the section: it says that, if we have an inessential vertex type, splitting a worrisome petal does not impede another worrisome petal from being split.
Lemma 4.19.
Let $\tau$ be an inessential vertex type in $\mathrm{Wh}_{\Gamma}$. Then, there exists a unique essential vertex type $\widehat{\tau}>\tau$ obtained by splitting only the worrisome petals in $\tau$.
Proof.
We will start by constructing $\widehat{\tau}$. We proceed in the following way: we take each based partition $\tau_{i}$ of $\tau$, and we split each worrisome petal in $\tau_{i}$ as much as possible, until there is no worrisome petal left. By Lemma 4.18, the previous procedure is independent of the order in which we split the petals, so we will always get to the vertex type $\widehat{\tau}$, which is essential, since it has no worrisome petal. ∎
Definition 4.20 (Essential cover).
We define the essential cover of an inessential vertex type $\tau$ as the unique essential vertex type $\widehat{\tau}$ obtained by splitting only the worrisome petals in $\tau$.
Lemma 4.21.
Let $\mathcal{A}$ be an inessential collection of compatible canonical outer-automorphisms. Then the complex $\mathrm{Peripheral}(\mathcal{A})$ is contractible.
Proof.
Let $\tau=\tau(\mathcal{A})$, and let $\mathfrak{t}>\tau$ be a vertex type obtained from $\tau$ by splitting at least one worrisome petal, and only worrisome petals. Let $\mathrm{Wh}_{\geq\mathfrak{t}}$ be the subposet of $\mathrm{Wh}_{\Gamma}$ consisting of vertex types greater than or equal to $\mathfrak{t}$. We have that the associated subcomplexes $|\mathrm{Wh}_{\geq\mathfrak{t}}|$ cover $\mathrm{Peripheral}(\mathcal{A})$, since in every vertex type in $\mathrm{Peripheral}(\mathcal{A})$ at least one of the worrisome petals in $\tau$ has been split (recall the remark after Definition 4.16). We have that each $|\mathrm{Wh}_{\geq\mathfrak{t}}|$ is contractible, since it is a cone over $\mathfrak{t}$. Even more, if we have two different $\mathfrak{t}_{1}$, $\mathfrak{t}_{2}$, by Lemma 4.18, they must be compatible; so, by Lemma 5.6 in [1], they have a least upper bound $\mathfrak{t}_{3}$ and it is clear that $|\mathrm{Wh}_{\geq\mathfrak{t}_{1}}|\cap|\mathrm{Wh}_{\geq\mathfrak{t}_{2}}|=|\mathrm{Wh}_{\geq\mathfrak{t}_{3}}|$. This means that we have covered $\mathrm{Peripheral}(\mathcal{A})$ by contractible complexes, whose intersection is also contractible. As a consequence we can apply the Quillen-McCord theorem (or Quillen fiber lemma) and deduce that $\mathrm{Peripheral}(\mathcal{A})$ is homotopic to the nerve of the covering.
[…]
Since $\mathrm{Peripheral}(\mathcal{A})\subseteq\mathcal{C}(\mathcal{A})$ and both are contractible we have:
Lemma 4.22.
The relative homology groups $H_{i}(\mathcal{C}(\mathcal{A}),\mathrm{Peripheral}(\mathcal{A}))$ vanish for all $i\geq 0$.
5. Computing $H^{*}(\Sigma\mathrm{POut}(A_{\Gamma}))$
We want to compute the cohomology groups $H^{q}(\Sigma\mathrm{POut}(A_{\Gamma}))$ using the equivariant spectral sequence associated to the action of $\Sigma\mathrm{POut}(A_{\Gamma})$ on $\mathrm{MM}_{\Gamma}$. Recall that if $G$ is a group acting simplicially on a contractible complex $X$, the first page of this equivariant spectral sequence (see, for example, Proposition 7.3 in Chapter VII of [4]) is
(1) $E_{1}^{p,q}=\prod_{\begin{subarray}{c}\sigma\in\left|W_{\Gamma}\right|,\\ \operatorname{dim}(\sigma)=p\end{subarray}}H^{q}(\operatorname{Stab}(\sigma))$
The differentials on the $E_{1}$ page are a combination of restriction and coboundary maps. We note that $\mathcal{B}(\tau)$, the canonical basis for the stabilizers of vertices, allows us to give a concrete description of the entries on the first page of this spectral sequence.
We know that the stabilizer of a cell $\sigma=\tau^{0}<\cdots<\tau^{k}$ is the stabilizer of $\tau^{0}$, so we can consider what happens in the case of a vertex $\tau$. The stabilizers are free abelian, so $H_{1}(\operatorname{Stab}(\tau),\mathbb{Z})$ is free abelian with generating set $\left\{\left[\bar{C}_{I}^{j}\right]\mid\bar{C}_{I}^{j}\in\mathcal{B}(\tau)\right\}$. As a consequence, $H^{1}(\operatorname{Stab}(\tau))$ is generated by the dual basis $\left\{\left[\bar{C}_{I}^{j}\right]^{*}\mid\bar{C}_{I}^{j}\in\mathcal{B}(\tau)\right\}$, we will denote $\bar{\gamma}_{I}^{j}:=\left[\bar{C}_{I}^{j}\right]^{*}$.
For free abelian groups, $H^{*}(\mathbb{Z}^{n})$ is an exterior algebra generated by one-dimensional classes, so the set $\left\{\bar{\gamma}_{I}^{j}\mid\bar{C}_{I}^{j}\in\mathcal{B}(\tau)\right\}$ is a generating set for the cohomology of $\mathrm{Stab}(\sigma)$. As a consequence, the set of all products of $q$ distinct $\bar{\gamma}_{I}^{j}$ classes is a generating set for $H^{q}(\operatorname{Stab}(\sigma))$. Therefore, we may choose as generating set for $H^{q}(\operatorname{Stab}(\sigma))$ the collection of dual classes of all subsets of canonical outer-automorphisms $\mathcal{A}\subset\mathcal{B}(\tau)$ with $|\mathcal{A}|=q$. Taking all of this into account, we see that in $E_{1}^{p,q}$ we have a basis element for each compatible collection $\mathcal{A}$ of canonical outer-automorphisms with $|\mathcal{A}|=q$ and each $\sigma$ of dimension $p$ where $\mathcal{A}\subseteq\mathcal{B}(\sigma)$.
5.1. The essential filtration
We aim to prove that the spectral sequence collapses at $E_{2}$. To do that we will construct a filtration through essential and inessential vertex types. We begin with a few useful remarks. Let $G\curvearrowright X$ be a group action, with $X$ a contractible simplicial complex and $X/G=W$ a finite fundamental domain. Let $F_{\bullet}(X)$ be the chain complex of $X$ and for any cell $\sigma\in X$ consider the stabilizer $\mathrm{Stab}(\sigma)$ and the induced module $\mathbb{Z}\uparrow^{G}_{\mathrm{Stab}(\sigma)}$. Then
$F_{p}(X)=\bigoplus_{\begin{subarray}{c}\sigma\in W\\ |\sigma|=p\end{subarray}}\mathbb{Z}\uparrow^{G}_{\mathrm{Stab}(\sigma)}.$
Now, for the homology equivariant spectral sequence we have:
$E^{1}_{p,q}=H_{q}(G,F_{p}(X))=H_{q}(G,\bigoplus_{\begin{subarray}{c}\sigma\in W\\ |\sigma|=p\end{subarray}}\mathbb{Z}\uparrow^{G}_{\mathrm{Stab}(\sigma)})=\bigoplus_{\begin{subarray}{c}\sigma\in W\\ |\sigma|=p\end{subarray}}H_{q}(G,\mathbb{Z}\uparrow^{G}_{\mathrm{Stab}(\sigma)})=\bigoplus_{\begin{subarray}{c}\sigma\in W\\ |\sigma|=p\end{subarray}}H_{q}(\mathrm{Stab}(\sigma),\mathbb{Z}).$
In a similar way, taking into account that $W$ is finite
$E^{p,q}_{1}=H^{q}(G,\textrm{Hom}_{\mathbb{Z}}(F_{p}(X),\mathbb{Z}))=H^{q}(G,\bigoplus_{\begin{subarray}{c}\sigma\in W\\ |\sigma|=p\end{subarray}}\textrm{ Coind }^{G}_{\mathrm{Stab}(\sigma)}\ \mathbb{Z})=\bigoplus_{\begin{subarray}{c}\sigma\in W\\ |\sigma|=p\end{subarray}}H^{q}(G,\mathbb{Z}\uparrow^{G}_{\mathrm{Stab}(\sigma)})=\bigoplus_{\begin{subarray}{c}\sigma\in W\\ |\sigma|=p\end{subarray}}H^{q}(\mathrm{Stab}(\sigma)).$
Since in our case all stabilizers are free abelian, we have that
$H^{q}(\mathrm{Stab}(\sigma))=\textrm{Hom}_{\mathbb{Z}}\left(H_{q}(\mathrm{Stab}(\sigma)),\mathbb{Z}\right).$
As a consequence, $E_{1}^{\bullet,q}$ can be obtained by dualizing $E^{1}_{\bullet,q}$, so if $E^{1}_{\bullet,q}$ has homology zero always except in degree zero where it is free abelian of rank $k$, the same will happen for $E_{1}^{\bullet,q}$. Therefore, for simplicity, we will work with the homology spectral sequence. We want to construct a filtration to compute the homology of $E^{1}_{\bullet,q}$. Because of Lemma 4.19, we know that for each inessential vertex $\tau$ we can find its essential cover $\widehat{\tau}$, an essential vertex in $\mathrm{Peripheral}(\mathcal{A})$.
Definition 5.1 ($s(\mathcal{A})$).
Let $\mathcal{A}$ be a family of compatible canonical outer-automorphisms. If $\mathcal{A}$ is essential we put $s(\mathcal{A})=0$, in other case let $\tau=\tau(\mathcal{A})$ and $\widehat{\tau}$ its essential cover, then $s(\mathcal{A})$ is the distance between $\tau$ and $\widehat{\tau}$, i.e. the length of any unrefinable chain $\tau<\tau^{1}<\cdots<\tau^{s(\mathcal{A})}=\widehat{\tau}$ .
We claim that there is a finite filtration
$F^{0}_{\bullet}\subset F^{1}_{\bullet}\subset\cdots\subset E^{1}_{\bullet,q}$
where $F_{p}^{s}=\bigoplus\mathbb{Z}_{\mathcal{A}}$, and the sum ranges over the families $\mathcal{A}\subset\mathcal{B}(\sigma)$ with $\sigma$ a $p$-cell such that $s(\mathcal{A})\leq s$ and $|\mathcal{A}|=q$. Moreover
$F_{\bullet}^{s}/F^{s-1}_{\bullet}=\bigoplus_{\mathcal{A}}\textrm{ Chain complex of the pair }(\mathcal{C}(\mathcal{A}),\mathrm{Peripheral}(\mathcal{A}))$
ranging over the set of compatible families $\mathcal{A}$, with $|\mathcal{A}|=q$ and $s(\mathcal{A})=s$; and
$F^{0}_{\bullet}=\bigoplus_{\mathcal{A}}\textrm{ Chain complex of }\mathrm{supp}(\mathcal{A})$
where now the sum is over the set of compatible families $\mathcal{A}$ such that $|\mathcal{A}|=q$ and $\mathcal{A}$ essential. To see it we have to check that the differential map $d^{1}$ of $E^{1}_{p,q}$ maps $F_{\bullet}^{s}$ to $F_{\bullet}^{s}$. We do this in our next lemma.
Lemma 5.2.
There is a finite filtration
$F^{0}_{\bullet}\subset F^{1}_{\bullet}\subset\cdots\subset E^{1}_{\bullet,q}$
with
$F_{p}^{s}=\bigoplus\{\mathbb{Z}_{\mathcal{A}}\ |\ \mathcal{A}\subseteq\mathcal{B}(\sigma),\sigma\in|\mathrm{Wh}_{\Gamma}|\textrm{ a }p\textrm{-cell},|\mathcal{A}|=q,s(\mathcal{A})\leq s\}$
such that $F_{\bullet}^{s}/F^{s-1}_{\bullet}$ is the sum of the chain complexes of pairs $(\mathcal{C}(\mathcal{A}),\mathrm{Peripheral}(\mathcal{A}))$ and $F^{0}_{\bullet}$ is the sum of chain complexes of $\mathrm{supp}(\mathcal{A})$ for $\mathcal{A}$ essential.
Proof.
The differential
$d^{1}:E^{1}_{p,q}\rightarrow E^{1}_{p-1,q}$
comes from
$d^{1}_{\sigma}:H_{q}(\mathrm{Stab}(\sigma),\mathbb{Z})\rightarrow\bigoplus H_{q}(\mathrm{Stab}(\sigma_{i}),\mathbb{Z})$
where $\sigma=\tau^{0}<\cdots<\tau^{p}$ and $\sigma_{i}$ is the result of removing $\tau^{i}$ from $\sigma$. So we have two possible cases:
  • If $i\neq 0$: $\mathrm{Stab}(\sigma_{i})=\mathrm{Stab}(\sigma)=\mathrm{Stab}(\tau^{0})$.
  • If $i=0$: $\mathrm{Stab}(\sigma_{0})=\mathrm{Stab}(\tau^{1})$.
As a consequence, for $\mu_{\mathcal{A}}\in F_{p}^{s}$ corresponding to the $p$-cell $\sigma$ we have:
$d^{1}_{\sigma}(\mu_{\mathcal{A}})=\sum_{i=0}^{p}(-1)^{i}cores_{i}(\mu_{\mathcal{A}})$
where corescores denote the corestriction map $H_{q}(\mathrm{Stab}(\sigma))\xrightarrow{cores}H_{q}(\mathrm{Stab}(\sigma_{i}))$. So $cores_{i}(\mu_{\mathcal{A}})\in F^{s}_{p}$ for $i\in\{1,...,p\}$, since they correspond to the same copy of $\mathcal{A}$, i.e., they are in $\mathcal{C}(\mathcal{A})$. In the case when $i=0$ we distinguish two cases:
Case 1: If we have obtained $\tau^{1}$ from $\tau^{0}$ without splitting any worrisome petal, then $\mathcal{A}\subset\mathcal{B}(\tau^{1})$, so $\tau^{1}\in\mathcal{C}(\mathcal{A})$ and the image is $F^{s}_{p}$.
Case 2: If we have obtained $\tau^{1}$ from $\tau^{0}$ by splitting a worrisome petal $I$ into $I_{1}$ and $I_{2}$:
Let $S=\mathrm{Stab}(\tau),S_{1}=\mathrm{Stab}(\tau^{1})$, so $S<S_{1}$. Let $\mathcal{A}\subseteq\mathcal{B}(\tau)$ such that $\bar{C}_{I}^{i}\in\mathcal{A}$, and $\mathcal{A}_{1}$, $\mathcal{A}_{2}$ be the sets formed by all the elements in $\mathcal{A}$ except for $\bar{C}_{I}^{i}$ which is replaced by $\bar{C}_{I_{1}}^{i}$ and $\bar{C}_{I_{2}}^{i}$ respectively. So $|\mathcal{A}_{1}|=|\mathcal{A}_{2}|=q$ and
$cores(\mu_{\mathcal{A}})\in\mathbb{Z}\textrm{-span}\left\{\mu_{\mathcal{A}_{j}}\ |\ j\in\{1,2\}\right\}.$
We take $\tau(\mathcal{A})$ and $\tau(\mathcal{A}_{1})$. All the based partitions are equal except for $\tau(\mathcal{A})_{i}$ and $\tau(\mathcal{A}_{1})_{i}$, which differ in two petals:
$\tau(\mathcal{A})_{i}=\{\{i\},Q_{1},...,Q_{k},I,M\}$
and
$\tau(\mathcal{A}_{1})_{i}=\{\{i\},Q_{1},...,Q_{k},I_{1},M^{\prime}\}$
where $M^{\prime}=M\cup I_{2}$ and both $M$ and $M^{\prime}$ contain the minimal element in $\Gamma-\mathrm{st}(v_{i})$ (recall Definition 4.10). Since $M$ and $M^{\prime}$ contain the minimal element, they cannot be worrisome; so the only possible difference between worrisome petals is the one between $I$ and $I_{1}$ so $\tau(\mathcal{A})$ and $\tau(\mathcal{A}_{1})$ both have the same essential cover $\widehat{\tau}$. We know that $I$ can be split at least once more than $I_{1}$, so distance from $\tau(\mathcal{A})$ to $\widehat{\tau}$ is bigger than the one from $\tau(\mathcal{A}_{1})$ to $\widehat{\tau}$; therefore, $s(\mathcal{A}_{1})<s(\mathcal{A})$ and the image is in $F^{s}_{p}$. The case of $\tau(\mathcal{A}_{2})$ is symmetric. ∎
As a consequence, all the previous lemmas together with Lemma 4.22 imply that
$H_{i}(F_{\bullet}^{l}/F^{l-1}_{\bullet})=0$
for all $i$, and we have, using also Lemma 4.15:
Lemma 5.3.
$H_{i}(E^{1}_{\bullet,q})=H_{i}(F^{0}_{\bullet})=\begin{cases}\mathbb{Z}^{m},\ \ \ i=0\\ 0,\ \ \ \ \ \ i>0\end{cases}$
where $m$ is the number of essential compatible families $\mathcal{A}$ with $|\mathcal{A}|=q$.
5.2. The $E_{2}$ page
By Lemma 5.3, we get
Proposition 5.4.
The non-zero entries on the $E_{2}$ page of the spectral sequence are concentrated in the $(0,q)$-column, and the entries are free abelian groups of rank the number of essential compatible families $\mathcal{A}$ with $|\mathcal{A}|=q$.
By definition, every essential set of compatible canonical basis elements $\mathcal{A}$ is associated to an essential vertex type $\tau(\mathcal{A})$, so we do also have:
Corollary 5.5.
Let $K_{q}$ be the number of rank $q$ essential vertex types in $\mathrm{Wh}_{\Gamma}$, then
$H^{q}(\Sigma\mathrm{POut}(A_{\Gamma}))=\mathbb{Z}^{K_{q}}.$
Remark.
The number $K_{q}$ can be computed algorithmically (see the Appendix A).
Theorem 5.6.
The algebra $H^{*}(\Sigma\mathrm{POut}(A_{\Gamma}))$ is generated by the one-dimensional classes $\bar{\gamma}_{I}^{j}$ where the $\bar{C}_{I}^{j}$ are canonical outer-automorphisms.
Proof.
By Lemma 4.4, $\Sigma\mathrm{POut}(A_{\Gamma})$ is generated by $\left\{\bar{C}_{I}^{j}\mid C_{I}^{j}\text{ is a canonical automorphism}\right\}$, its first homology is generated by the associated classes $\left[\bar{C}_{I}^{j}\right]$ and $H^{1}\left(\Sigma\mathrm{POut}(A_{\Gamma})\right)$ is generated by the dual basis $\left\{\bar{\gamma}_{I}^{j}\right\}$. As we have seen in Proposition 5.4, the $E_{2}$ page is concentrated in a single column and the groups in $E_{2}^{0,q}$ are free abelian. The generating set corresponds to products of elements of the form $\bar{\gamma}_{K}^{j}=\sum_{I\in K}\bar{\gamma}_{I}^{j}$. These cohomology classes $\bar{\gamma}_{K}^{j}$ come from a generating set for the stabilizer of an essential vertex type, so the partial conjugations $C_{I}^{j}$ are canonical automorphisms. ∎
Theorem A follows from Corollary 5.5 and Theorem 5.6.
Example 5.7.
We will compute the cohomology groups $H^{q}(\Sigma\mathrm{POut}({A_{\Gamma}}))$ for
$A_{\Gamma}=\langle v_{1},v_{2},v_{3},v_{4},v_{5}\ |\ [v_{1},v_{2}]=1\rangle$
by computing the number of essential vertices in $\mathrm{Wh}_{\Gamma}$.
$\Gamma=$$v_{1}$$v_{2}$$v_{4}$$v_{3}$$v_{5}$
Figure 2. The graph $\Gamma$
As we can see, $A_{\Gamma}$} is nearly a free group, as a consequence, $\mathrm{Wh}_{\Gamma}$ will have some similarities with $\mathrm{Wh}_{4}$ (a representation of $\mathrm{Wh}_{4}$ can be found at Figure 2 in [13]); however, the adjacency between $v_{1}$ and $v_{2}$ results in their associated based partitions always being compatible, so the number of possible combinations increases.
The subposet of $\mathrm{Wh}_{\Gamma}$ associated to vertex types where based partitions associated to $v_{1}$ and $v_{2}$ are trivial is the same as the subposet of $\mathrm{Wh}_{4}$ where the based partitions associated to one of the vertices are always trivial, i.e, nine rank 1 and nine rank 2 vertex types.
The subposet of $\mathrm{Wh}_{\Gamma}$ associated to vertex types where the only non-trivial based partitions are the ones associated to $v_{1}$ is the same as the subposet of $\mathrm{Wh}_{4}$ where the only non-trivial based partitions are the ones associated to a single vertex, i.e., three rank 1 vertex types and 1 rank 2 vertex type. The same occurs with the subposet associated to vertex types where the only non-trivial based partitions are the ones associated to $v_{2}$. Since $v_{1}$ and $v_{2}$ are adjacent, the subposet of $\mathrm{Wh}_{\Gamma}$ associated to vertex types where the only non-trivial based partitions are the ones associated to $v_{1}$ and $v_{2}$ is the direct product of their respective subposets. This direct product results in six rank 1, eleven rank 2, six rank 3 and one rank 4 vertex types.
We are left with the possible combinations of $\{v_{1},v_{2}\}$ with $\{v_{3},v_{3},v_{5}\}$. There are precisely two rank 2 vertex types with non-trivial based partitions at $v_{1}$ and $v_{3}$: if we split $v_{4}$ in one based partition we have to split $v_{5}$ in the other one, and vice versa. The same happens for any pair $(v_{i},v_{j})$ with $i\in\{1,2\}$ and $j\in\{3,4,5\}$. As a consequence, we have twelve rank 2 vertex types associated to these pairs. There will also be two rank 3 vertices associated to the triple $(v_{1},v_{2},v_{3})$: if we have split $v_{4}$ in the based partition associated to $v_{3}$, then we can split $v_{5}$ in the based partitions associated to both $v_{1}$ and $v_{2}$ and the other way around. In the same way, we have two rank 3 vertices associated to the triples $(v_{1},v_{2},v_{4})$ and $(v_{1},v_{2},v_{5})$.
Once we described all the possible vertex types, we have to see which of them are essential. For each $v_{i}$, we have that $\Gamma-\mathrm{st}(v_{i})$ has three connected components, this means that a vertex type $\tau$ will be inessential if and only if it has a based partition $\tau_{i}=\{\{v_{i}\},\{v_{j}\},\{v_{k},v_{l}\}\}$ such that $v_{j}$ is the minimal element in $\Gamma-\mathrm{st}(v_{i})$ and the petal $\{v_{k},v_{l}\}$ can be split. As a consequence, between the fifteen rank 1 vertex types, ten of them are essential (the ones where the minimal element has not been split).
For the rank 2 vertex types, the nine associated to $\{v_{3},v_{4},v_{5}\}$ are essential, since they are maximal (any other split would produce a crossing). Between the eleven associated to $\{v_{1},v_{2}\}$, six of them are essential; the two vertex types where one of the based partitions is totally split and the rest are trivial, and the ones where the minimal element has not been split in any of the based partitions. All the twelve rank 2 vertex types associated to mixed pairs $(v_{i},v_{j})$ with $i\in\{1,2\}$ and $j\in\{3,4,5\}$ are essential, since the only based partition which admits a split is the one associated to $v_{2}$ in case $i=1$ or vice versa, and that based partition is trivial.
For the rank 3 vertex types, the six associated to triples $(v_{1},v_{2},v_{j})$ with $j\in\{3,4,5\}$ are essential since they are maximal. It remains to check the six vertex types associated to the pair $(v_{1},v_{2})$. In these vertex types one of the based partitions is totally split and the other one will be $\tau_{i}=\{\{v_{i}\},\{v_{j}\},\{v_{k},v_{l}\}\}$. The essential ones are the ones where $v_{j}$ is not the minimal element, so there are four of them.
In summary, we have the next information:
Rank Vertex types Essential
0 1 1
1 15 10
2 32 27
3 12 10
4 1 1
Table 1. $\Gamma$-Whitehead poset
And, as a consequence
[…]
Observe that even if $A_{\Gamma}$} looks close to $F_{4}$, the differences in the homology of the symmetric outer-automorphisms group are remarkable: by the formula given in Lemma 5.2 in [9], we get
[…]
In Appendix A, we give a brief explanation of an algorithm that computes not just the number of essential vertex types in $\mathrm{Wh}_{\Gamma}$, but also the $n$-simplices in $|\mathrm{Wh}_{\Gamma}|$ and, can be used to compute the $E^{1}$ page of the equivariant spectral sequence for the $\Sigma\mathrm{POut}(A_{\Gamma})$ action on $MM_{\Gamma}$. The algorithm can be implemented in Python and be used to check the previous example; though we encourage the inquisitive reader to try to do it by hand (a couple of paper sheets will be probably needed).
6. Computing $H^{*}(\Sigma\mathrm{PAut}(A_{\Gamma}))$
6.1. Some previous results
We begin with some well known results about homology and cohomology of right-angled Artin groups, which we will be necessary later, and follow from the fact that the Salvetti complex associated to a graph $\Gamma$, is a classifying space for the RAAG $A_{\Gamma}$ (see [11]).
Proposition 6.1.
Let $R$ be a commutative ring and $A_{\Gamma}$ a RAAG, then
$H_{k}(A_{\Gamma},R)\cong R^{N_{k}}$
where $N_{k}$ is the number of $k$-cliques in $\Gamma$. The cohomology group $H^{k}(A_{\Gamma},R)$ has the same rank.
Proposition 6.2.
Let $R$ be a commutative ring and $A_{\Gamma}$ a RAAG. Let $n$ be the number of vertices of $\Gamma$ and $\{e^{1},...,e^{n}\}$ the standard basis of $H^{1}({A_{\Gamma}},R)$. Then the cohomology ring $H^{\ast}(A_{\Gamma},R)$ is the exterior algebra $\bigwedge\left(R\right)^{n}$ modulo the relations $e^{i}\wedge e^{j}=0$ whenever the vertices in $\Gamma$ associated to $i$ and $j$ are not adjacent.
Note that, in particular, the previous proposition implies that $H^{\ast}(A_{\Gamma},R)$ is generated in degree 1.
6.2. Applying the Leray-Hirsch Theorem
We are assuming that $Z({A_{\Gamma}})$ is trivial (see the remark after Theorem 2.5), so we have a short exact sequence:
$1\to A_{\Gamma}\to\Sigma\mathrm{PAut}(A_{\Gamma})\to\Sigma\mathrm{POut}(A_{\Gamma})\to 1$
[…]
Exactly as in Section 6 in [9], there is a fibration which allows us to apply the Leray-Hirsch Theorem. This theorem implies that the differential $d^{2}$ of the spectral sequence is trivial, so the sequence collapses at page 2. This second page is given by
$E_{2}^{*,q}=H^{*}(\Sigma\mathrm{POut}(A_{\Gamma}),H^{q}(A_{\Gamma}))=H^{*}(\Sigma\mathrm{POut}(A_{\Gamma}),\mathbb{Z}^{t_{q}})$
where $t_{q}$ is the number of $q$-cliques in $\Gamma$. Following the argument of [9], we only need to check $H^{k}({A_{\Gamma}})$ is $\mathbb{Z}$-free, which is given by Proposition 6.1 and the statement of the next lemma.
Lemma 6.3.
The map
$\iota^{*}:H^{*}(\Sigma\mathrm{PAut}(A_{\Gamma}))\rightarrow H^{*}(A_{\Gamma})$
is surjective, where $\iota:A_{\Gamma}\rightarrow\Sigma\mathrm{PAut}(A_{\Gamma})$ is the inclusion.
Proof.
We claim first that the map
$H^{1}(\Sigma\mathrm{PAut}(A_{\Gamma}))\rightarrow H^{1}({A_{\Gamma}})$
is surjective. To see it, note that the injection $A_{\Gamma}\hookrightarrow\Sigma\mathrm{PAut}(A_{\Gamma})$ sends each basis element $v_{j}$ to the symmetric automorphism
$C_{[n]\backslash\{\mathrm{st}(v_{j})\}}^{j}=C_{I_{1}}^{j}\cdot C_{I_{2}}^{j}\cdots C_{I_{k}}^{j}$
where the $I_{1},...,I_{k}$ are connected components in $\Gamma-\mathrm{st}(v_{j})$. $H_{1}\left(\Sigma\mathrm{PAut}(A_{\Gamma})\right)$ and $H_{1}\left({A_{\Gamma}}\right)$ are free abelian groups with the same generating set as $\Sigma\mathrm{PAut}(A_{\Gamma})$ and $A_{\Gamma}$ respectively. The partial conjugations form a generating set for $\Sigma\mathrm{PAut}(A_{\Gamma})$ that gives rise to a minimal generating set for the abelianization, so the map
$H_{1}\left({A_{\Gamma}}\right)\rightarrow H_{1}\left(\Sigma\mathrm{PAut}(A_{\Gamma})\right)$
is injective and the claim follows. As, by Proposition 6.2, $H^{*}({A_{\Gamma}})$ is generated in degree one, we see that the map
$H^{*}(\Sigma\mathrm{PAut}(A_{\Gamma}))\rightarrow H^{*}({A_{\Gamma}})$
is also surjective.
Corollary 6.4.
Let $K_{i}$ be the number of k essential vertex types of rank $i$ (generators of $H^{i}(\Sigma\mathrm{POut}(A_{\Gamma}))$) and $N_{j}$ the number of $j$-cliques in $\Gamma$ (generators of $H^{j}({A_{\Gamma}})$, with $N_{0}=1$). Then
$H^{q}(\Sigma\mathrm{PAut}(A_{\Gamma}))\cong\bigoplus_{i+j=q}H^{i}(\Sigma\mathrm{POut}(A_{\Gamma}))\otimes H^{j}({A_{\Gamma}})$
$\cong\bigoplus_{i+j=q}H^{i}(\Sigma\mathrm{POut}(A_{\Gamma}),H^{j}(A_{\Gamma}))\cong\bigoplus_{i+j=q}\mathbb{Z}^{K_{i}\cdot N_{j}}$
Proof.
The result follows directly from Corollary 5.5, Proposition 6.1 and the Leray-Hirsch Theorem. ∎
So, with the same argument as in Corollary 6.4 in [9], we get:
Corollary 6.5.
The ring $H^{*}(\Sigma\mathrm{PAut}(A_{\Gamma}))$ is generated in degree 1.
This concludes the proof of Theorem B.
7. The cohomology ring $H^{*}(\Sigma\mathrm{PAut}({A_{\Gamma}}))$
In this section we will give a ring presentation that we conjecture gives the cohomology ring $H^{*}(\Sigma\mathrm{PAut}({A_{\Gamma}}))$. We also prove some partial results that provide some evidence for this conjecture. Let $i,j$ be two vertices which are not adjacent in $\Gamma$, let $D^{j},D^{i}$ be the respective dominant components of $\Gamma-\mathrm{st}(v_{i})$, $\Gamma-\mathrm{st}(v_{j})$ and $\mathcal{C}$ a shared component. As in Section 5, we denote by $\gamma_{A}^{i}$ the dual class $C_{A}^{i}]^{*}$. We want to prove that the next two relations hold in $H^{*}(\Sigma\mathrm{PAut}({A_{\Gamma}}))$:
  • (1)
    $\gamma_{D^{j}}^{i}\wedge\gamma_{D^{i}}^{j}=0$
  • (2)
    $\gamma_{\mathcal{C}}^{i}\wedge\gamma_{\mathcal{C}}^{j}=\gamma_{\mathcal{C}}^{i}\wedge\gamma_{D^{i}}^{j}+\gamma_{D^{j}}^{i}\wedge\gamma_{\mathcal{C}}^{j}$
Proposition 7.1.
With the previous notation, relation $(1)$ holds in $H^{*}(\Sigma\mathrm{PAut}(A_{\Gamma}))$.
Proof.
We will prove the relation through an epimorphism from $\Sigma\mathrm{PAut}(A_{\Gamma})$ to $F_{2}$, the free group generated by $\{x,y\}$. We define
$\varphi_{1}:\Sigma\mathrm{PAut}(A_{\Gamma})\rightarrow F_{2}$
induced by
$C_{D^{j}}^{i}\mapsto x,\quad C_{D^{i}}^{j}\mapsto y,\quad\alpha\mapsto 1\textrm{ for }\alpha\textrm{ any other generator. }$
Taking into account the presentation of $\Sigma\mathrm{PAut}({A_{\Gamma}})$ given in Theorem iii), we see that there is no possible commutation between $C_{D^{j}}^{i}$ and $C_{D^{i}}^{j}$ so this map extends to a well defined group homomorphism. The map $\varphi_{1}$ induces a homomorphism $\overline{\varphi}_{1}^{*}$ between the first cohomology groups:
$\overline{\varphi}_{1}^{*}:H^{1}(F_{2})\rightarrow H^{1}(\Sigma\mathrm{PAut}(A_{\Gamma})).$
With a suitable order in the generators of $\Sigma\mathrm{PAut}(A_{\Gamma})$, $\overline{\varphi}_{1}$ can be represented by the following matrix:
$\begin{pmatrix}1&0&0&\cdots&0\\ 0&1&0&\cdots&0\end{pmatrix}$
so, dualizing, $\overline{\varphi}_{1}^{*}$ corresponds to the transpose, which means $\overline{\varphi}_{1}^{*}(x^{*})=\gamma_{D^{j}}^{i}$ and $\overline{\varphi}_{1}^{*}(y^{*})=\gamma_{D^{i}}^{j}$. In $H^{1}(F_{2})$, we have that $x^{*}\wedge y^{*}=0$; and the naturality of the cup product with respect to group homomorphisms implies
$\varphi_{1}^{*}(x^{*}\wedge y^{*})=\varphi_{1}^{*}(x^{*})\wedge\varphi_{1}^{*}(y^{*})=\gamma_{D^{j}}^{i}\wedge\gamma_{D^{i}}^{j}=0.$
Proposition 7.2.
With the notation established before, relation $(2)$ also holds in $H^{*}(\Sigma\mathrm{PAut}(A_{\Gamma}))$.
Proof.
We argue as in Proposition 7.1. We define
$\varphi_{2}:\Sigma\mathrm{PAut}(A_{\Gamma})\rightarrow F_{2}$
induced by
$C_{D^{j}}^{i}\mapsto x,\quad C_{D^{i}}^{j}\mapsto y,\quad C_{\mathcal{C}}^{i}\mapsto x^{-1},\quad C_{\mathcal{C}}^{j}\mapsto y^{-1},\quad\alpha\mapsto 1\textrm{ for }\alpha\textrm{ any other generator. }$
We will verify again that this extends to a well defined group homomorphism. Taking into account the presentation given in Theorem iii), we see that the only relations that involve the preimages of $x^{\pm 1}$ and $y^{\pm 1}$ are $C_{\mathcal{C}}^{i}C_{D^{j}}^{i},C_{\mathcal{C}}^{j}]=1$ and $C_{\mathcal{C}}^{j}C_{D^{i}}^{j},C_{\mathcal{C}}^{i}]=1$. It is enough to check that $\varphi_{2}(C_{\mathcal{C}}^{i}C_{D^{j}}^{i}C_{\mathcal{C}}^{j})=\varphi_{2}(C_{\mathcal{C}}^{j}C_{\mathcal{C}}^{i}C_{D^{j}}^{i})$ (the other relation will follow by symmetry):
$\varphi_{2}(C_{\mathcal{C}}^{i}C_{D^{j}}^{i}C_{\mathcal{C}}^{j})=x^{-1}xy^{-1}=y^{-1}=y^{-1}x^{-1}x=\varphi_{2}(C_{\mathcal{C}}^{j}C_{\mathcal{C}}^{i}C_{D^{j}}^{i}).$
Again, with a suitable order in the generators of $\Sigma\mathrm{PAut}(A_{\Gamma})$, $\overline{\varphi}_{2}$ can be represented by the following matrix:
$\begin{pmatrix}1&0&-1&0&0&\cdots&0\\ 0&1&0&-1&0&\cdots&0\end{pmatrix}$
so $\overline{\varphi}_{2}^{*}$, the induced homomorphism between the first cohomology groups, corresponds to the transpose, and we get
$\overline{\varphi}_{2}^{*}(x^{*})=\gamma_{D^{j}}^{i}-\gamma_{\mathcal{C}}^{i}\ \ \textrm{ and }\ \ \overline{\varphi}_{2}^{*}(y^{*})=\gamma_{D^{i}}^{j}-\gamma_{\mathcal{C}}^{j}.$
By the properties of the cup product
$\overline{\varphi}_{2}^{*}(x^{*}\wedge y^{*})=\overline{\varphi}_{2}^{*}(x^{*})\wedge\overline{\varphi}_{2}^{*}(y^{*})=(\gamma_{D^{j}}^{i}-\gamma_{\mathcal{C}}^{i})\wedge(\gamma_{D^{i}}^{j}-\gamma_{\mathcal{C}}^{j})=0$
and reordering and applying Proposition 7.1, we obtain
$\gamma_{\mathcal{C}}^{i}\wedge\gamma_{\mathcal{C}}^{j}=\gamma_{\mathcal{C}}^{i}\wedge\gamma_{D^{i}}^{j}+\gamma_{D^{j}}^{i}\wedge\gamma_{\mathcal{C}}^{j}.$
Definition 7.3.
We define $R$ as the graded commutative ring generated by all the $\gamma_{A}^{i}$, where $1\leq i\leq n$ and $A$ ranges over the connected components in $\Gamma-\mathrm{st}(v_{i})$, with the relations
  • (1)
    $\gamma_{A}^{i}\wedge\gamma_{A}^{i}=0$
  • (2)
    $\gamma_{A}^{i}\wedge\gamma_{B}^{j}=-\gamma_{B}^{j}\wedge\gamma_{A}^{i}$
  • (3)
    $\gamma_{D^{j}}^{i}\wedge\gamma_{D^{i}}^{j}=0$
  • (4)
    $\gamma_{\mathcal{C}}^{i}\wedge\gamma_{\mathcal{C}}^{j}=\gamma_{\mathcal{C}}^{i}\wedge\gamma_{D^{i}}^{j}+\gamma_{D^{j}}^{i}\wedge\gamma_{\mathcal{C}}^{j}$ for each shared component $\mathcal{C}$ of $v_{i},v_{j}$ .
We will denote by $R_{q}$ the term of degree $q$ in R.
By Propositions 7.1 and 7.2, we have:
Corollary 7.4.
Let $R$ be the ring of Definition 7.3. Then there exists a ring epimorphism
$\phi:H^{*}(\Sigma\mathrm{PAut}(A_{\Gamma}))\twoheadrightarrow R.$
so
$\textrm{dim}(R_{q})\geq\textrm{dim}(H^{q}(\Sigma\mathrm{PAut}(A_{\Gamma})).$
We want to prove now that $\phi$ is an isomorphism in degree 2. To do that, we will construct an epimorphism from the standard $\mathbb{Z}$-basis of
(2) $H^{2}(\Sigma\mathrm{PAut}(A_{\Gamma}))\cong\bigoplus_{i+j=2}\mathbb{Z}^{K_{i}\cdot N_{j}}\cong\mathbb{Z}^{K_{2}\cdot N_{0}}\oplus\mathbb{Z}^{K_{1}\cdot N_{1}}\oplus\mathbb{Z}^{K_{0}\cdot N_{2}}$
to a $\mathbb{Z}$-basis of $R_{2}$. In the formula, we have $N_{0}=1$, and $N_{1}$, $N_{2}$ are the number of vertices and edges in $\Gamma$ respectively.
Definition 7.5.
We will denote by $\mathcal{B}_{1}$ the standard $\mathbb{Z}$-basis of $H^{2}(\Sigma\mathrm{PAut}({A_{\Gamma}}))$, which is bijective to the union of the following sets:
  • Type 1 elements: the set of all rank 2 essential vertex types.
  • Type 2 elements: the set of tall pairs $(\tau,j)$ where $\tau$ is a rank 1 essential vertex and $v_{j}$ a vertex in $\Gamma$.
  • Type 3 elements: the set of all edges in $\Gamma$.
Definition 7.6.
Let $\mathcal{B}_{2}$ be the family of elements the of $R_{2}$ consisting of those $\gamma_{A}^{i}\gamma_{B}^{j}$ such that $A\subset\Gamma-\mathrm{st}(v_{i}),B\subset\Gamma-\mathrm{st}(v_{j})$ are connected components and when $v_{i}\notin\mathrm{st}(v_{j})$
  • $A,B$ cannot be shared and equal.
  • $A,B$ cannot be both the respective dominant components.
Lemma 7.7.
$\mathcal{B}_{2}$ is a $\mathbb{Z}$-basis for $R_{2}$.
Proof.
$\mathcal{B}_{2}$ generates all $R_{2}$, since by relation $(3)$ $\gamma_{A}^{i}\gamma_{B}^{j}=0$ if $A,B$ are both the respective dominant components, and if $A=B=\mathcal{C}$ then by relation $(4)$ $\gamma_{\mathcal{C}}^{i}\gamma_{C}^{j}$ can be obtained as a $\mathbb{Z}$-linear combination of $\gamma_{\mathcal{C}}^{i}\gamma_{D^{i}}^{j}$ and $\gamma_{D^{j}}^{i}\gamma_{C}^{j}$. $\mathcal{B}_{2}$ is clearly linearly independent since there are no other relations left to make a linear combination zero if the coefficients are not all zero. ∎
Remark.
If we have two compatible based partitions $\tau_{i},\tau_{j}$ and a petal $P\in\tau_{i}$ does not contain the dominant component $D^{j}$, then $P$ can always be split in the case when it is formed by more than one connected component, and the new based partition $\tau_{i}^{\prime}$ will still be compatible with $\tau_{j}$. Recall as well that if $\Gamma-\mathrm{st}(v_{i})$ and $\Gamma-\mathrm{st}(v_{j})$ have different minimal elements, then one of the minimal elements must lie in the respective dominant component.
Proposition 7.8.
There exists an injective map $\varphi:\mathcal{B}_{1}\rightarrow\mathcal{B}_{2}$.
Proof.
We are going to construct $\varphi$ explicitly for each of the types of elements of $\mathcal{B}_{1}$ considering several possible cases. First, we will fix the image of elements of type 1.
Type 1: Rank 2 vertex types $\tau$.
1.1: $\tau$ has a single non-trivial based partition $\tau_{i}=\{\{v_{i}\},P_{1},P_{2},P_{3}\}$: We may assume that the minimal element of $\Gamma-\mathrm{st}(v_{i})$ is in $P_{1}$. Then, since $\tau$ is essential, $P_{2}$ and $P_{3}$ cannot admit a split, so they must consist of a unique connected component, because otherwise they could be split, since all of the other based partitions are trivial. We set
$\varphi(\tau)=\gamma^{i}_{P_{2}}\gamma^{i}_{P_{3}}\in\mathcal{B}_{2}.$
1.2: $\tau$ has two non-trivial based partitions $\tau_{i}=\{\{v_{i}\},P_{1},P_{2}\}$ and $\tau_{j}=\{\{v_{j}\},Q_{1},Q_{2}\}$: We may assume that the respective minimal elements are in $P_{1}$ and $Q_{1}$. We will see that when $v_{i}\notin\mathrm{lk}(v_{j})$, at least one of the $P_{1},Q_{1}$ contains the respective dominant component. In other case, the minimal element is the same in both $\tau_{i}$ and $\tau_{j}$, since in order to be different, the minimal associated to $v_{i}$ should be in $\mathrm{st}(v_{j})$ (or the other way around) and it would be in the dominant component. But then we would have $P_{1}\cap Q_{1}\neq\emptyset$, and this would be a crossing since the dominant components are not in those petals, which means that the partitions would not compatible, so we have a contradiction. It is obvious that $P_{2}$ and $Q_{2}$ cannot be shared and equal, because in that case we would have a crossing too. We distinguish two subcases.
1.2.1: $v_{i},v_{j}$ are not linked, $P_{1}$ contains the dominant component $D^{j}$, and for the other dominant component $D^{i}$ we have $D^{i}\subsetneq Q_{2}$. Then $P_{2}\subset Q_{2}$ and we set
$\varphi(\tau)=\gamma^{i}_{P_{2}}\gamma^{j}_{D^{i}}\in\mathcal{B}_{2}$
1.2.2: In other case, we set:
$\varphi(\tau)=\gamma^{i}_{P_{2}}\gamma^{j}_{Q_{2}}\in\mathcal{B}_{2}$
Let us check that the claim in 1.2.1. We know that $P_{2}$ and $Q_{2}$ cannot be split. Since $P_{2}$ does not contain the dominant, then it must consist of a unique connected component. On the other hand, $Q_{2}=\{D^{i}\cup\mathcal{C}_{1}\cup\cdots\cup\mathcal{C}_{k}\}$, where each $\mathcal{C}_{i}$ is shared, because subordinate components do not produce crossings and can always be split. But these $\mathcal{C}_{1},...,\mathcal{C}_{k}$ must be in $P_{2}$, to impede the split of $Q_{2}$ being valid. As a consequence, $P_{2}=\mathcal{C}$ and $Q_{2}=D^{i}\cup\mathcal{C}$. Since $P_{2}$ consists of a unique connected component which is not $D^{j}$ we have that $\gamma^{i}_{P_{2}}\gamma^{j}\in\mathcal{B}_{2}$.
In case 1.2.2, we have already seen that $P_{2}$ and $Q_{2}$ cannot contain both the respective dominant components. So it remains to check that both $P_{2}$, $Q_{2}$ consist of a unique connected component and that $P_{2}\neq Q_{2}$, i.e., they are not shared and equal. We have three possibilities
  • $v_{i}\in\mathrm{st}(v_{j})$. In this case there cannot be crossings between $\tau_{i}$ and $\tau_{j}$, so any split is allowed. As a consequence, since they are essential, both $P_{2}$ and $Q_{2}$ must be formed by a unique connected component; since they are adjacent there are no shared components so $\gamma^{i}_{P_{2}}\gamma^{j}_{Q_{2}}\in\mathcal{B}_{2}$.
  • $v_{i}\notin\mathrm{st}(v_{j})$ and both $P_{1}$ and $Q_{1}$ contain the respective dominant components. Since $P_{2}$ and $Q_{2}$ cannot admit a split, both $P_{2}$ and $Q_{2}$ must be formed by a unique connected component; they can not be equal since we would have a crossing so $\gamma^{i}_{P_{2}}\gamma^{j}_{Q_{2}}\in\mathcal{B}_{2}$.
  • $v_{i}\notin\mathrm{st}(v_{j})$ and $P_{1}$ contains the dominant component and $Q_{2}$ is just the dominant component. By the same argument as before, we know that $P_{2}$ must be formed by a unique connected component. $Q_{2}$ is also formed by just one connected component and it is not shared so $\gamma_{P_{2}}^{j}\gamma_{Q_{2}}^{i}\in\mathcal{B}_{2}$.
It is clear that the map $\varphi$ constructed so far is injective. Besides, notice that for all of the elements of the form $\gamma^{i}_{U_{1}}\gamma^{i}_{U_{2}}$ or $\gamma^{i}_{U_{1}}\gamma^{j}_{U_{2}}$ that we have in the image, the $U_{1},U_{2}$ never contain the respective minimal element.
Type 2: Pairs $(\tau,j)$ for $\tau$ a rank 1 essential vertex and $v_{j}$ a vertex in $\Gamma$. In $\tau$ every based partition is trivial except for one $\tau_{i}=\{\{v_{i}\},P_{1},P_{2}\}$. We will assume that the minimal element is in $P_{1}$ and $P_{2}$ must consist of a unique connected component; in other case, it would admit a split. Again, we consider several subcases.
2.1: Assume $v_{i}\notin\mathrm{st}(v_{j})$, the minimal elements in $\Gamma-\mathrm{st}(v_{j})$ and $\Gamma-\mathrm{st}(v_{i})$ are different and the one associated to $v_{j}$ lies in a shared component $\mathcal{C}$, so the one associated to $v_{i}$ must lie in $D^{j}$:
2.1.1: For a pair $(\tau,j)$ with $\tau_{i}=\{\{v_{i}\},P_{1},\mathcal{C}\}$, we set:
$\varphi(\tau,j)=\gamma^{i}_{D^{j}}\gamma^{j}_{\mathcal{C}}$
2.1.2: For a pair $(\tau^{\prime},i)$ with $\tau^{\prime}_{j}=\{\{v_{j}\},Q_{1},D^{i}\}$, we set:
$\varphi(\tau^{\prime},i)=\gamma^{j}_{D^{i}}\gamma^{i}_{\mathcal{C}}$
2.2: Assume $v_{i}\notin\mathrm{st}(v_{j})$, the minimal element in $\Gamma-\mathrm{st}(v_{j})$ is in the dominant component $D^{i}$, the minimal element in $\Gamma-\mathrm{st}(v_{i})$ is in a subordinate component $\mathcal{S}$ and the only non-trivial based partition in $\tau$ is $\tau_{i}=\{\{v_{i}\},P_{1},D^{j}\}$. We set:
$\varphi(\tau,j)=\gamma^{i}_{\mathcal{S}}\gamma^{j}_{D^{i}}$
2.3: In any other case, for the pair $(\tau,j)$ with $\tau_{i}=\{\{v_{i}\},P_{1},P_{2}\}$ such that the minimal element in $\Gamma-\mathrm{st}(v_{i})$ lies in $P_{1}$ and the minimal element in $\Gamma-\mathrm{st}(v_{j})$ lies in $U$, we set:
$\varphi(\tau,j)=\gamma^{i}_{P_{2}}\gamma^{j}_{U}$
Type 3: Edges $(v_{i},v_{j})\in E(\Gamma)$. Consider the connected components $U_{i}$ of $\Gamma-\mathrm{st}(v_{i})$ and $U_{j}$ of $\Gamma-\mathrm{st}(v_{j})$ that contain the minimal elements and ser
$\varphi(i,j)=\gamma^{i}_{U_{i}}\gamma^{j}_{U_{j}}$
During the proof we have already checked that the elements chosen in the image of $\varphi$ lie in $\mathcal{B}_{2}$; equivalently, that they are of the form $\gamma^{i}_{A}\gamma^{j}_{B}$ where $\neq B, AA$ and $B$ consist of a unique connected component in $\Gamma-\mathrm{st}(v_{i})$ and $\Gamma-\mathrm{st}(v_{j})$ respectively, and $A,B$ are not the respective dominant components at the same time. It remains to prove that the map is injective. Looking at the construction of $\varphi$ we see that in each case we get elements $\gamma^{i}_{A}\gamma^{j}_{B}$ such that:
  • Type 1: neither $A$ nor $B$ contain the respective minimal element.
  • Type 2: in cases 2.1.1 and 2.2, both $A$ and $B$ contain the minimal element and $i\notin\mathrm{st}(j)$. In case 2.3, $B$ contains the minimal element and $A$ does not. In case 2.1.2 neither $A$ nor $B$ contain the respective minimal element.
  • Type 3: both $A$ and $B$ contain the minimal element and $i\in\mathrm{st}(j)$.
As a consequence, the possible sources of problems for injectivity are:
Type 1 and 2.1.2: We have to check that for $v_{i},v_{j}$ as in case 2.1.2, there is no rank 2 vertex type $T$ such that:
$\varphi(T)=\gamma^{j}_{D^{i}}\gamma^{i}_{\mathcal{C}}.$
In that case we have that $T_{i}=\{\{v_{i}\},P_{1},P_{2}\}$ and $T_{j}=\{\{v_{j}\},Q_{1},Q_{2}\}$, where $P_{1}$ and $Q_{1}$ contain the minimal elements, so $\mathcal{C}\subseteq Q_{1}$. If $\varphi(T)=\gamma^{j}_{D^{i}}\gamma^{i}_{\mathcal{C}}$, then $D^{i}\subseteq Q_{2}$ and $\mathcal{C}=P_{2}$, but then there is crossing since $Q_{1}\cap P_{2}=\mathcal{C}$ and neither $Q_{1}$ nor $P_{2}$ contain the respective dominant component, so we have a contradiction.
2.1.1, 2.2 and type 3: Images will always be different between type 2 and type 3 elements because in cases 2.1.1 and 2.2, $i\notin\mathrm{st}(v_{j})$ and in type 3 $(v_{i},v_{j})$ is an edge. Images in cases 2.1.1 and 2.2 will also be different since in the first case $A,B$ are a dominant and a shared component and in the second one is dominant and the other subordinate.
Pairs $(\tau,i),(\tau^{\prime},j)$: Let $\tau$ be a vertex type such that the only non-trivial based partition is $\tau_{i}=\{\{v_{i}\},P_{1},P_{2}\}$ and $\tau^{\prime}$ such that the only non-trivial based partition is $\tau_{j}=\{\{v_{j}\},Q_{1},Q_{2}\}$. Then, it could happen that the images of those pairs are $\gamma^{i}_{A}\gamma^{j}_{B}$ in the first case and $\gamma^{j}_{B}\gamma^{i}_{A}$ in the second one, breaking injectivity. Case 2.1 has been divided into 2.1.1 and 2.1.2 precisely to avoid it. Case 2.2 is not problematic since in 2.1 subordinate components do not appear and in 2.3 only one of the components contain the minimal element. Case 2.3 is also not problematic, since if we have $(\tau,j)$ and $(\tau^{\prime},i)$ such that
$\varphi(\tau,j)=\gamma^{i}_{A}\gamma^{j}_{B}$
and
$\varphi(\tau^{\prime},i)=\gamma^{j}_{B}\gamma^{i}_{A}$
then the argument in case 2.3 implies that $A$ contains and does not contain the minimal element, which is a contradiction. ∎
So we have
Theorem 7.9.
The map $\phi$ of Corollary 7.4 is an isomorphism in degrees 1 and 2, so, in particular
$H^{2}(\Sigma\mathrm{PAut}(A_{\Gamma}))\cong R_{2}.$
Proof.
The fact that $\phi$ is surjective implies that $\textrm{dim}(R_{2})\geq\textrm{dim}(H^{2}(\Sigma\mathrm{PAut}(A_{\Gamma}))$ and Proposition 7.8 implies that both dimensions are equal. ∎
Conjecture 7.10 (Generalized Brownstein-Lee Conjecture).
The map $\phi$ of Corollary 7.4 is an isomorphism.
8. The Torelli subgroup
Let $G$ be a group, the Torelli subgroup $\mathrm{IA}(G)$ is the kernel of the homomorphism:
$\phi:\mathrm{Aut}(G)\rightarrow\mathrm{Aut}(G^{ab})\cong\mathrm{Aut}(H_{1}(G)).$
In general, not much is known about these groups. In the case of a RAAG $A_{\Gamma}$}, Day found in [6] a finite generating system for $\mathrm{IA}({A_{\Gamma}})$, and Wade proved in his PhD thesis [17] that $H_{1}(\mathrm{IA}({A_{\Gamma}}))$ is the free abelian group on that set of generators. Following the idea of Corollary 6.9 in [9] (pointed by Fred Cohen), we can obtain some information about $H^{*}(\mathrm{IA}({A_{\Gamma}}))$ from Theorem B.
Corollary 8.1.
The injection $\Sigma\mathrm{PAut}({A_{\Gamma}})\hookrightarrow\mathrm{IA}({A_{\Gamma}})$ induces a split epimorphism
$H^{*}(\mathrm{IA}({A_{\Gamma}}))\twoheadrightarrow H^{*}(\Sigma\mathrm{PAut}(A_{\Gamma})).$
Furthermore, the suspension of $B\Sigma\mathrm{PAut}({A_{\Gamma}})$ is homotopic to a bouquet of spheres, and it is a retract of the suspension of $B\mathrm{IA}({A_{\Gamma}})$.
Proof.
Once it has been proven that $H^{*}(\Sigma\mathrm{PAut}({A_{\Gamma}}))$ is generated by the one-dimensional classes (Corollary 6.5), and that $H_{1}(\mathrm{IA}({A_{\Gamma}}))$ is free abelian (Theorem 4.23 in [17]), the result follows by the same arguments as in Corollary 6.9 in [9]. ∎
Appendix A Algorithm to compute the $K_{q}$:
The main theorems of this paper (Theorem A and B) give us a formula to compute the cohomology groups $H^{q}(\Sigma\mathrm{POut}({A_{\Gamma}}))$ and $H^{q}(\Sigma\mathrm{POut}({A_{\Gamma}}))$. In the case when $A_{\Gamma}$} is the free group $F_{n}$, it is possible to give a closed formula in terms of $n$ only (Lemma 5.2 in [9]). However, that cannot be done for arbitrary RAAGs; we can, nonetheless, implement an algorithm that can be used for explicit computations.
The idea is simple: for each vertex $v_{i}\in\Gamma$, we take all the connected components in $\Gamma-\mathrm{st}(v_{i})$ and combine them in all possible ways to obtain $\mathcal{P}_{i}$, the set of the based partitions associated to each $v_{i}$. Then, we consider the cartesian product of the $\mathcal{P}_{i}$’s. We have to remove from the cartesian product the elements that are not vertex types, so for each SIL-pair $(v,w)$ we check which based partitions cross and we remove all the elements which have those two based partitions at the same time.
Once we have all the vertex types, we select those of rank $q$ and check if there is some based partition which does not contain the minimal element and admits a split; if there is none, then the vertex type is essential.
A code implemented for Python which computes the Whitehead poset, the number of essential vertices for each rank, the number of simplices in $|\mathrm{Wh}_{\Gamma}|$ and the $E^{1}$ page for the $\Sigma\mathrm{POut}({A_{\Gamma}})$ action on $\mathrm{MM}_{\Gamma}$ can be found at https://github.com/peioardaiz-gale/Codes-for-the-generalized-Whitehead-poset/blob/main/whitehead-poset_essential-vertices_E1-page.py.
Algorithm 1 Algorithm to compute $K_{q}$, the number of essential vertex types of rank $q$
1:for $1\leq q\leq n-2$ do
2:  $K_{q}\leftarrow 0$
3:end for
4:for $1\leq i\leq n$ do
5:  Compute a list $\mathcal{P}_{i}$ with all the valid based partitions of $\Gamma-\mathrm{st}(v_{i})$
6:end for
7:for $\tau\in\mathcal{T}$ do
8:  if the based partitions in $\tau\in\mathcal{T}$ are pairwise compatible then
9:   $\tau\in\mathrm{Wh}_{\Gamma}$
10:  end if
11:end for
[…]
13:  $K_{q}\leftarrow 0$
14:  for $\tau\in\mathrm{Wh}_{\Gamma}$ do
15:   if $rk(\tau)=q$ then
16:     if $\tau$ is essential then
17:      $K_{q}++$
18:     end if
19:   end if
20:  end for
21:end for